### v1

# On the Hyperparameter Loss Landscapes of Machine Learning Algorithms

Abstract
Despite the recent success in a plethora of hyperparameter optimization (HPO) methods for machine learning (ML) models, the intricate interplay between model hyperparameters (HPs) and predictive losses (a.k.a fitness), which is a key prerequisite for understanding HPO, remain notably underexplored in our community. This results in limited explainability in the HPO process, rendering a lack of human trust and difficulties in pinpointing algorithm bottlenecks. In this paper, we aim to shed light on this black box by conducting large-scale fitness landscape analysis (FLA) on $1,500$ HP loss landscapes of $6$ ML models with more than $11$M model configurations, across $67$ datasets and different levels of fidelities. We reveal the first unified, comprehensive portrait of their topographies in terms of smoothness, neutrality and modality. We also show that such properties are highly transferable across datasets and fidelities, providing fundamental evidence for the success of multi-fidelity and transfer learning methods. These findings are made possible by developing a dedicated FLA framework that incorporates a combination of visual and quantitative measures. We further demonstrate the potential of this framework by analyzing the NAS-Bench-101 landscape, and we believe it is able to faciliate fundamental understanding of a broader range of AutoML tasks.
1 Introduction
In the past decade, considerable efforts have been invested in developing hyperparameter optimization (HPO) techniques to automate the laborious task of hyperparameter (HP) tunning for machine learning (ML) models. Many successful approaches (Bergstra et al., 2011; Snoek et al., 2012; Hutter et al., 2011; Srinivas et al., 2010; Karnin et al., 2013; Li et al., 2017; Falkner et al., 2018; Awad et al., 2021) have significantly advanced this field, and they have been empirically shown to outperform both manual configuration (Hutter et al., 2019; Bischl et al., 2023; Santu et al., 2022; Yang & Shami, 2020) and random search (Bergstra & Bengio, 2012).
HPO is often casted as a black-box optimization problem (BBOP), where the goal is to search for an HP configuration $\bm{\lambda}\in\bm{\Lambda}=\Lambda_{1}\times\dots\Lambda_{n}$ with an objective value $\mathcal{L}(\bm{\lambda})$ as small as possible without any explicit knowledge of the ML loss function $\mathcal{L}:\bm{\Lambda}\to\mathbb{R}$. Existing methods (see examples above) to this end essentailly comprises $3$ key components: $i)$: a search space, $ii)$: an optimization strategy, and $iii)$: model evaluation. While the development of both efficient searching mechanisms and evaluation strategies has received considerable attention in recent years, the intricate interplay between model HPs and predictive losses, which plays a pivotal role in understanding HPO problems, remain notably underexplored. Such lack of knowledge in turn hampers the transparency and explainability (Dwivedi et al., 2023) of HPO solvers, which often function as black-boxes as well. Consequently, this results in limited human trust in HPO methods and hinders their wide-spread application (Drozdal et al., 2020; Bischl et al., 2023) Unfortunately, given the high-dimensional, hybrid nature of HP configuration space, it is far from trivial to open up such black box.
The fitness landscape metaphor, which was pioneered by Wright in 1932 in evolutionary biology, has been widely recognized as a powerful tool for analyzing BBOPs in the evolutionary computation community (Malan, 2021). It can be envisioned as a (hyper-)surface as formed by objective values, over the high-dimensional space of possible configurations (Romero et al., 2013). Since the 90s, a plethora of fitness landscape analysis (FLA) methods have been developed to conduct exploratory analysis on landscape characteristics of BBOPs (Zou et al., 2022). Such methods are useful, as they are able to extract landscape measures that are indicative of problem difficulty and how a certain search mechanism would perform on it (Smith-Miles & Lopes, 2012; Hutter et al., 2014b; Qasem & Prügel-Bennett, 2010). This knowledge would then advance the understanding of the problem characteristics (Huang & Li, 2023), assist the selection and configuration of problem solvers (Kerschke et al., 2019; Schede et al., 2022), navigate the design of new algorithms (Qasem & Prügel-Bennett, 2010), and enhance the explainability and trust for optimization (Thomson et al., 2023).
Figure 1: $2$D visualization of HP loss landscapes for: (a-d) CNN on $9$ HPs and $6,480$ configurations, (e-f) XGBoost regressor on $5$ HPs and $14,960$ configurations, under different scenarios. Colors indicate ranks of configurations (lower values are better). Coordinates are projected using UMAP.
Recently, the use of FLA in analyzing HP and the related AutoML loss landscapes has also received considerable attention. Various works have studied diverse structural characteristics of these landscapes including neutrality, modality, and fitness distance correlation (e.g., Pushak & Hoos (2022); Teixeira & Pappa (2022); Pimenta et al. (2020); Schneider et al. (2022)). However, such works suffer from limited setups and fail to interrogate the connection between landscape characteristics and the success of HP optimizers, which often run in a wide range of scenarios (e.g., different models, datasets and fidelities). It remains unclear whether the HP loss landscapes induced on different settings share certain characteristics or patterns, and how such commonalities could be potentially exploited by HP optimizers. On the other hand, we argue that existing analytical methods are insufficient to provide a comprehensive analysis and comparision on HP loss landscapes since:
  • The ability to visualize landscapes is crucial for enabling intuitive understandings of their complex topologies (Michalak, 2019). However, HP loss landscapes are notoriously difficult to visualize in a human-comprehensible fashion due to their high-dimensional nature. Some existing methods address this problem by plotting only one or two HPs each time (e.g., Friedman (2001); Akiba et al. (2019)), which fail to provide an integrated view of the global landscape structure. Other works applied dimensionality reduction techniques to project the landscape into $2$D space (e.g., Michalak (2019); Biedenkapp et al. (2018); Walter et al. (2022)), but the resulting plot is not able to preserve the overall topography as well as neighborhood structure of the landscape.
  • There is no tangible method for quantifying the similarity between different HP loss landscapes. Despite general FLA metrics could already capture informative landscape characteristics, practices in automated algorithm selection demonstrate that domain-specific metrics are also crucial as a complementary source of information for better characterizing the target problem (Smith-Miles, 2008; Smith-Miles & Lopes, 2012). However, none of the prior works have considered such perspectives when comparing HP loss landscapes.
The overarching goal of this paper is to gain an integral view of the HP loss landscapes induced on different scenarios and thereby provide new insights to the community. To this end, we develop a dedicated landscape analysis framework to enable comprehensive analysis and comparisions among HP loss landscapes. It incorporates 1: a novel neighborhood-aware HP loss landscape visualization method applicable to high-dimensions, 2: a series of FLA metrics quantifying landscape structural characteristics, and 3: $3$ similarity metrics that leverage rankings of HP configurations to allow for informative landscape similarity quantification in the HPO context. Through empirical analysis on $1,500$ landscapes across $6$ ML models and $67$ datasets with more than $11$ million configurations, we are ambitious to advance the understanding of the following four fundamental HPO scenarios:
HP Landscapes of Test Loss Versus Training Loss. ‘Overfitting’ is one of the biggest interests and concerns in the ML community (Ng, 1997; Caruana et al., 2000; Recht et al., 2019; Belkin et al., 2018; Roelofs et al., 2019; Ishida et al., 2020). However, there is a lack of in-depth understanding on how test loss correlates with training loss across a broad HP landscape, and what specific properties distinguish regions that generalize well from poorly generalized ones. In this paper, by using our fitness landscape analysis framework, we find that the test loss landscapes resemble their training counterparts in terms of both structural characteristics and performance rankings (see, e.g., Figure 1 (a) versus (b)), and configurations with small training error are likely to achieve a mild generalization error. However, significant discrepancies can also occur (see, e.g., Figure 1 (e) versus (f)) depending on both the choice of certain HP combinations and the dataset at hand. In such cases, struggling to reduce the training loss has little or even negative effect to refining the generalization loss.
HP Loss Landscapes Across Fidelities. Given the time-demanding nature of model evaluation, multi-fidelity HPO methods (Karnin et al., 2013; Kandasamy et al., 2016; Li et al., 2017; Kandasamy et al., 2017; Falkner et al., 2018; Awad et al., 2021) have achieved prominent performance by more efficient resource allocation. However, the validity of their underpinned assumption, i.e., the ranking of configuration performance would stay close to the ground truth under fidelities (Bischl et al., 2023), remains unclear (Pushak & Hoos, 2022). Our empirical results are highly inspiring to support such assumption and show that landscapes with lower fidelities are highly consistent with full-fidelity landscapes w.r.t. both structural characteristics and performance ranks (Figure 1 (c)).
HP Loss Landscapes Across Datasets. Leveraging priors obtained from previous tasks to expedite the learning process for new tasks is another promising direction for HPO (Feurer et al., 2015b; Bardenet et al., 2013; Wistuba et al., 2015b; Kim et al., 2017; Rakotoarison et al., 2022; Wistuba et al., 2015c; Vanschoren, 2018; Swersky et al., 2013). These approaches are grounded in the hypothesis that ‘knowledge’ about HPO landscapes—such as configuration quality, hyperparameter importance and their interactions (Hutter et al., 2014a; Watanabe et al., 2023b; Probst et al., 2019; van Rijn & Hutter, 2017)—is correlated across related datasets when defined under a certain distance measure. A natural question that arises is whether this knowledge remains consistent when applied to a broader range of tasks. Our results on a diverse set of $67$ datasets show that performance rankings, HP importance and their interactions, are largely shared across tasks (Figure 1 (d)).
HP Loss Landscapes Across Models. Methods rooted in Bayesian optimization (e.g., Snoek et al. (2012); Bergstra et al. (2011); Feurer et al. (2015a); Kandasamy et al. (2017)) and search space pruning (e.g., Wistuba et al. (2015a); Perrone & Shen (2019); Wang et al. (2020)) implicitly assume that the quality of configurations is locally correlated rather than a full randomness. This is also true for meta-heuristic methods (e.g., Friedrichs & Igel (2005); Lessmann et al. (2005); Cawley (2001); Guo et al. (2008)), which are even more dependent on specific landscape properties. While it may seem intuitive that HP loss landscapes would differ depending on the target ML model, in practice the fact is often that common HPO methods perform robustly for different models. This implies that, despite superficial differences, the general family of HP loss landscapes may share certain inherent patterns/properties. We verified this hypothesis by synthesizing the results from diverse FLA metrics characterizing HP loss landscape geometry combined with visual inspections (see, e.g., Figure 1 (a, e)). The results gathered from $1,500$ landscapes of $6$ ML models under different scenarios, reveal a universal picture of the HP loss landscapes. In this picture, HP loss landscapes are smooth, nearly unimodal, containing a large number of neutral regions; configurations with similar performance are locally clustered; the landscape becomes flatter around the optimum configuration.
2 HPO Landscape Analysis Methods
This section introduces our analytical framework developed to enable exploratory analysis on different HP loss landscapes and perform comparisons between them. Due to the page limit, we only provide a brief sketch of these methods here while more detailed discussions are in Appendix B.
HP Loss Landscape Construction. The HP loss landscape can be formulated as a triplet $\langle\bm{\Lambda},\mathcal{L},\mathcal{N}\rangle$ with three ingredients: $i)$ a search space $\bm{\Lambda}$ of feasible configurations that consists of pre-evaluated, discretized grid values (see Appendix F.1), $ii)$ a ML loss function $\mathcal{L}:\bm{\lambda}\to\mathbb{R}$, and $iii)$ a neighborhood structure $\mathcal{N}$ that specifies which configurations are neighbors to each other. Note that the form of $\mathcal{N}$ depends on a distance function $\mathrm{d}:\bm{\lambda}\times\bm{\lambda}\to\mathbb{N}$. Following Pushak & Hoos (2022), we define all categorical values of a HP to be distance $1$ from each other (i.e., the values are non-ordinal). For a numerical HP, we define the distance between two values to be the number of steps between them on the grid used for discretization. Such distance measure is able to mimic the tunning strategy of human experts when combined with elaborately designed grid values. Based on this, the total distance between two configurations $\bm{\lambda}_{i}$ and $\bm{\lambda}_{j}$ is then sum of the distances between the respective pairs of HP values, and we say they are neighbors to each other (i.e., $\bm{\lambda}_{j}\in\mathcal{N}(\mathcal{\bm{\lambda}}_{i})$), if $\mathrm{d}(\bm{\lambda}_{j},\bm{\lambda}_{i})=1$. Finally, the HPO landscape is constructed as a directed graph where the vertices are HP configurations and an improving edge $e_{i,j}\in\mathcal{E}$ is traced from $\bm{\lambda}_{i}$ to $\bm{\lambda}_{j}$ if $\bm{\lambda}_{j}\in\mathcal{N}(\bm{\lambda}_{i})$ and $\mathcal{L}(\bm{\lambda}_{j})<\mathcal{L}(\bm{\lambda}_{i})$. We say that a configuration $\bm{\lambda}_{\ell}$ is a local optimum if $\forall\bm{\lambda}^{\prime}\in\mathcal{N}(\bm{\lambda}_{\ell})$, we have $\mathcal{L}(\bm{\lambda}_{\ell})<\bm{\lambda}^{\prime}$. In addition, we say that $\bm{\lambda}_{j}$ is a neutral neighbor of $\bm{\lambda}_{i}$ if their performance difference is negligible ($\leq 1$‰).
Landscape Visualization. We develop a first-of-its-kind, highly interpretable method for visualizing the topography of high-dimensional HP loss landscapes by leveraging graph representation learning (Hamilton, 2020) combined with dimensionality reduction (Draganov et al., 2023) techniques. Specifically, we first extracted low-dimensional features for each node in the graph. To this end, we use HOPE (Ou et al., 2016) node embedding method because it could preserve high-order proximities between configurations. Then, we compress the obtained feature vectors into $2$ components using the UMAP (McInnes & Healy, 2018) algorithm, and thus allowing configurations to be laid out in $2$D scatter plot. To further refine the interpretability of the obtained plots, we additionally apply a linear interpolation and thereby generate a smooth landscape surface.
Table 1: Summary of the FLA metrics used in our landscape analysis framework.
Metrics
Symbol Domain What a Higher Value Implies
[…]
$\mathcal{L}$-ast
$[-1,1]$
HP Configurations with similar $\mathcal{L}$ values are more likely to be neighbors to each other.
[…]
$\rho_{a}$
$[-1,1]$
The landscape is smoother
Neutrality Distance Correlation
NDC
$[-1,1]$
The landscape is more likely to be flatter near the optimum.
[…]
$\bar{\nu}$
$[0,1]$
There are many ‘plateaus’ in the landscape.
No. Local Optima
$n_{\text{lo}}$
$\mathbb{N}^{+}$
There are many ‘valleys’ or ‘peaks’ in the landscape.
Mean Basin Size
$\bar{s}_{\mathcal{B}}$
$\mathbb{R}^{+}$
The local optima are hard to be escaped from.
 ${}^{1}$ewman (2010); ${}^{2}$einberger (1990), ${}^{3}$eidys & Stadler (2001)
Quantifying Landscape Characteristics. To quantitatively assess the structural characteristics of HP loss landscapes, we employ a series of dedicated FLA metrics summarized in Table 1 as surrogate features. There are many other metrics for characterizing landscape properties (see Zou et al. (2022) for a detailed review), but our selected ones are particularly useful for this study as they cover the most essential landscape properties (i.e., modality, neutrality and smoothness) that are related to algorithm behaviors. More importantly, they are intuitive enough even for non-experts in FLA.
Landscape Similarity in Terms of Performance Ranking. The comparison of rankings of HP configurations’ performance is the essence of a large corpora of HPO methods (Hutter et al., 2019). We thereby ground our landscape similarity measure of HP loss landscapes on the consistency of their performance ranks, denoted as $R(\mathcal{L}(\bm{\lambda}))$, to allow more informative results in the HPO context. Specifically, we use $3$ statistical metrics with complementary perspectives: 1)1)1 ) Spearman’s $\rho_{s}$, it measures the association of the performance ranks of configurations in two landscapes (Spearman, 1961), 2)2)2 ) Kaggle’s Shake-up metric (Trotman, 2019), it assesses the average movement of configuration rankings across two landscapes. 3)3)3 ) The $\gamma$-set similarity (Watanabe et al., 2023a), it quantifies the ratio of overlaps between top-$10\%$ regions of two landscapes divided by their unions.
In addition to these, to investigate the consistency of HP importance and interaction under different scenarios, We apply the widely used functional ANOVA method (Hutter et al., 2014a) to assess the variance contribution of every HP $\lambda\in\bm{\lambda}$ as well as their interactions up to the $3^{\text{rd}}$ order.
3 Experimental Setup
Table 2: Summarization of meta-information of our empirical study.
  Models Datasets Fidelities Summarization
 Model
Total
HPs
Total
Configs.
Cat.
Class.
Cat.
Reg.
Num.
Class.
Num.
Reg.
Image
Class.
Training
Data
Training
Epochs
Total
Configs.
# Land-
-scapes
XGB $5$ $14,960$ $15$ $7$ $17$ $18$$\{0.1,0.25,\text{all}\}$ $2.56$M $342$
RF $6$ $11,250$ $15$ $7$ $17$ $18$$\{0.1,0.25,\text{all}\}$ $1.92$M $342$
LGBM $5$ $13,440$ $15$ $7$ $17$ $18$$\{0.1,0.25,\text{all}\}$ $2.30$M $342$
DT $5$ $24,200$ $15$ $7$ $17$ $18$$\{0.1,0.25,\text{all}\}$ $4.14$M $342$
CNN $8$ $6,480$ $6$$\{0.1,0.25,\text{all}\}$$\{10,25,50\}$ $0.35$M $108$
[…]
  Total (Before accounting $5$-fold cross-validation): $\bm{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}11.15\textbf{M}}}$ $\bm{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}1,500}}$
Table 2 summarizes the meta-information of our empirical study, while detailed HP search space of each model and the principles we follow in designing them, are left in Appendix F.1. We first consider decision tree (DT(Safavian & Landgrebe, 1991) and three types of its ensembles: random forest (RF(Breiman, 2001), XGBoost (Chen & Guestrin, 2016) and LightGBM (Ke et al., 2017). We analyze the HP space of these models using the tabular benchmark proposed in Grinsztajn et al. (2022), which comprises $25$ regression and $32$ classification tasks (see Appendix F.2). These datasets span a broad range of complexities in terms of number of instances and features and are thus idea choice for comprehensive inspection of landscape characteristics. In addition to these, we also study convolutional neural networks (CNNs(Krizhevsky et al., 2012) on six classic image classification tasks (Appendix F.2) using a joint architecture and hyperparameter search (JAHS) (Bansal et al., 2022) space. We additionally consider another JAHS scenario, for which we adopt the NASBench-HPO (Klein & Hutter, 2019) data included in HPOBench (Eggensperger et al., 2021). This includes $62,208$ configurations of a feed-forward neural network (FCNet) evaluated on $4$ UCI datasets
For each dataset, unless predefined, we randomly split the data into training ($80\%$) and test ($20\%$) set. For all HP configurations $\bm{\lambda}\in\bm{\Lambda}$ of each model, we exhaustively evaluate $\mathcal{L}(\bm{\lambda})_{\text{train}}$ and $\mathcal{L}(\bm{\lambda})_{\text{test}}$ using $5$-fold cross-validation. Here, we use root mean squared error (RMSE) and $R^{2}$ score to serve as the loss function $\mathcal{L}$ for regression tasks, and for classification, we use accuracy and ROC-AUC. We control the fidelity of the training by varying the number of training instances to $\{10\%,25\%,100\%\}$ of the whole training data. For CNN, we additionally set the budget for the number of epochs to $\{10,25,50\}$ and thus obtain a total of $3\times 3$ different levels of fidelity. For FCNet, we vary fidelity by using meta-data at the $\{10,50,100\}$-th epoch. At the end, we obtain a total of $1,500$ landscapes with more than $11$M distinct HP configurations. To further demonstrate the transferability and potential impact of our proposed landscape analysis frameowork, we also employ it to analyze NASBench-101 (Ying et al., 2019), a well-known neural architecture search (NAS) benchmark.
4 Results and Analysis
In this section, we seek to investigate HP loss landscapes under the four scenarios posed in Section 1. We will start from providing an universal view of the general characteristics of HP loss landscapes of different ML models (Section 4.1). We then explore and compare landscapes under: $i)$ training and test setups (Section 4.2), $ii)$ different fidelities (Section 4.3), $iii)$ different datasets (Section 4.4).
Figure 2: 2D visualization of HP loss landscapes of $6$ ML models under different scenarios: (a) $\mathcal{L}_{\text{test}}$ landscape on baseline datasets (44059 for tree-based models, CIFAR-10 for CNN, protein structure for FCNet), (b) $\mathcal{L}_{\text{train}}$ landscape on baseline datasets, (c) Low-fidelity $\mathcal{L}_{\text{test}}$ landscape on baseline datasets, (d) $\mathcal{L}_{\text{test}}$ landscape on different datasets (44143 for tree-based models, Fashion MINIST for CNN, slice localization for FCNet). Colors indicate $R(\mathcal{L})$ (lower rank values are better).
Figure 3: Distribution of FLA metrics introduced in Table 1 for each model across all datasets for landscapes of 1)1)1 ) $\mathcal{L}_{\text{test}}$, 2)2)2 ) $\mathcal{L}_{\text{train}}$ and 3)3)3 ) $\mathcal{L}_{\text{test}_{\text{LF}}}$.
Figure 4: Distribution of Spearman, Shake-up and $\gamma$-set metrics between (a) $\mathcal{L}_{\text{test}}$ and $\mathcal{L}_{\text{train}}$, (b) $\mathcal{L}_{\text{test}}$ and $\mathcal{L}_{\text{test}_{\text{LF}}}$, (c) $\mathcal{L}_{\text{test}}$ across datasets. Medians are labeled beside each plot.
4.1 Overall Characteristics of HP Loss Landscape of ML Models
From landscape visualizations depicted in Figure 2 (a), we have a general impression that HP loss landscapes of ML models are highly structured and share certain patterns: they are relatively smooth; configurations are clustered in terms of performance; there is a highly distinguishable plateau consisting of prominent configurations, where the terrain becomes flatter. This impression is consistent with the FLA metrics reported in Figure 3, from which we see that landscapes for all models are:
Fairly smooth and clustered. The high $\mathcal{L}$-ast and $\rho_{a}$ values for $\mathcal{L}_{\text{test}}$ landscapes shown in Figure 3 (a) and (b) respectively imply that configurations with similar $\mathcal{L}_{\text{test}}(\bm{\lambda})$ tend to be locally connected, where a small change in $\bm{\lambda}$ is not likely to cause dramatic variation of $\mathcal{L}_{\text{test}}(\bm{\lambda})$. This observation is similar to the findings in reinforcement learning (RL) (Eimer et al., 2023), where the transitions between different parts of the HP landscapes of RL are also found to be quite smooth. This property makes the HP landscapes favorable to Bayesian optimization and search space pruning techniques, as it would be easier to separate the promising regions from the poorly performing ones. On the other hand, if the landscape is rugged instead, in which $\mathcal{L}_{\text{test}}(\bm{\lambda})$ of very different levels often mixed together, it would become more difficult for such techniques to identify a clear promising region.
Nearly unimodal. As shown in Figure 3 (e), we find that a considerable fraction of the $\mathcal{L}_{\text{test}}$ landscapes are unimodal, whereas for other landscapes, there could be a handful to dozens (DT) of local optima. This is somewhat contradictory to Pushak & Hoos (2022) at the first thought, in which the authors found that almost all landscapes they studied are nearly unimodal. However, when taking a closer look at the local optima in our landscapes, we find that they usually feature a small basin of attraction (Figure 3 (f)). This makes them relatively ‘shallow’, and thus would not pose significant obstacles to optimization.
However, beyond the results in Figure 3, we find that FCNet landscapes on the $4$ UCI datasets possess $24$ to $347$ local optima, with $\bar{s}_{\mathcal{B}}$ up to $2,372$ (Appendix D), implying a strong attraction for optimizers. Pushak & Hoos have also reported similar observations on these four landscapes, and they speculated the reason could be that these scenarios fall into the over-parameterized regime. While we agree with this reasoning, we seek to conduct further analysis on the local optima using the local optima network (LON) (Ochoa et al. (2008), Appendix B.3). We find that despite the pressence of many other local optima, the global optima still plays a pivotal role in the connectivity of the LON (Appendix D). Therefore, many local optima can eventually escape the global optimum via cetain strategies (e.g., a perturbation), though this may take additional efforts.
Highly neutral; planar around the optimum. As depicted in Figure 3 (d), we can clearly see that HP loss landscapes are often featured in high neutrality. This indicates that a large portion of $1$-bit moves in the landscape will result in subtle change in $\mathcal{L}_{\text{test}}(\bm{\lambda})$ (i.e., $\leq 1\text{\textperthousand}$). We postulate a major reason for this is the low effective dimensionality (Bergstra & Bengio, 2012) of HPO problems: usually only a small subset of all available HPs have obvious influence on performance. Despite landscape neutrality can largely vary with the choice on which HPs to analyze and their respective values, considering the fact that we have ready removed totally unimportant HPs from teh search space, moves with subtle performance shifts can actually be more prevalent than one may expect. Such phenomenon is more pronounced for the well-performing regions, as illustrated by the high NDC values in Figure 3 (c). It suggests that as we move closer to the global optimum, it is more likely to encounter neutral moves, and the landscape becomes flatter. This is in line with Probst & Boulesteix (2017); Pimenta et al. (2020) and practical experience: the gain of tuning HPs usually progressively decreases as approaching the best reachable performance. Such property also poses challenges to optimizers, as there is little gradient information that can be utilized for navigation towards fitter configurations (Muñoz et al., 2015).
Overall, despite certain exceptions, we see that the family of HP loss landscapes tend to share various high-level properties, whereas the exact topologies would vary with models. This explains why in practice, one can often expect an optimizer to work relatively robustly across a wide range of scenarios. In addition, most properties here also seem to generalize the NAS problems (Appendix C), except that we find the NAS landscapes tend to have lower neutrality and more local optima.
4.2 Training and Test HPO Landscapes
Figure 2 (a) and (b) provide a visual comparison of $\mathcal{L}_{\text{train}}$ and $\mathcal{L}_{\text{test}}$ landscapes for different models, respectively. We could infer from the plots that the structural characteristics of the $\mathcal{L}_{\text{train}}$ landscapes highly our previously discussed properties. On the other hand, for the performance rankings, we notice that $\mathcal{L}_{\text{train}}$ generally correlate with $\mathcal{L}_{\text{test}}$ for RF, DT, LGBM and CNN, whereas for XGBoost, there is significant shifts in performance between the two cases. We further quantitatively verify such observations using the FLA metrics and the landscape similarity metrics introduced in Section 2.
Structural characteristics. From Figure 3, we could clearly see that the landscape characteristics for $\mathcal{L}_{\text{train}}$ and $\mathcal{L}_{\text{test}}$ are highly consistent for most studied scenarios. More specifically, it is surprising to see that $\mathcal{L}_{\text{train}}$ landscapes tend to yield relatively higher $\mathcal{L}$-ast and $\rho_{a}$, suggesting a smoother and more structured terrain. Meanwhile, the NDC values are lower in the training scenarios. These observations imply that $\mathcal{L}_{\text{train}}$ landscapes are even more benign than $\mathcal{L}_{\text{test}}$ landscapes. In addition, we find that $\bar{\nu}$, $n_{\text{lo}}$ and $\bar{s}_{\mathcal{B}}$ values rarely change between $\mathcal{L}_{\text{train}}$ and $\mathcal{L}_{\text{test}}$ landscapes. Notably, local optima found in $\mathcal{L}_{\text{train}}$ and $\mathcal{L}_{\text{test}}$ landscapes are almost (if not all) identical. These indicate that the relative performance in local neighborhoods tend to remain similar in the two cases, despite the variations in their numerical values and the global performance rankings.
Landscape similarity in terms of performance rankings. We quantified the similarity between all pairs of $\mathcal{L}_{\text{train}}$ and $\mathcal{L}_{\text{test}}$ landscapes for all $5$ models using the three metrics introduced in Section 2 as shown in Figure 4 (a). Overall, we observe that $R(\mathcal{L}_{\text{train}})$ and $R(\mathcal{L}_{\text{test}})$ are globally correlated for all models except XGBoost, as indicated by the significant $\rho_{s}$ values (median $>0.7$) and low Shake-up metrics (median $<0.15$). However, when zooming into the top-$10\%$ regions, we find that the majority of our studied scenarios reveal low $\gamma$-set similarities. It indicates that the generalization gap is larger in prominent regions where configurations are highly adapted to the training set. This phenomenon is more severe for XGBoost, where the median $\gamma$-set similarity is only $0.07$, and there is also a poor $\rho_{s}$ value (median = $0.34$) and high Shake-up score (median = $0.25$).
In order to gain more insight into such generalization gaps for XGBoost, we create scatter plots of $\mathcal{L}_{\text{test}}$ versus $\mathcal{L}_{\text{train}}$ on dataset #44059 as shown in Figure 5 (a). We decompose the pattern into two modes: During the first mode, $\mathcal{L}_{\text{test}}$ highly correlates with $\mathcal{L}_{\text{train}}$ as it decreases, and the models in this stage underfit the data. In the next mode, as points struggle to further move on the $x$-axis ($\mathcal{L}_{\text{train}}$), they stagnate or even significantly increase on the $y$-axis ($\mathcal{L}_{\text{test}}$), indicating strong evidence of overfitting. In particular, we can see a plateauing trend near the $x$-axis, where some models overly excel on the training data, but performing poorly on the test set.
To further investigate which kinds of configurations are likely to lead to overfitting, we color the points with respect to their HP values as shown in Figure 5 (b-e). We are excited to see that the generated plots demonstrate clear patterns between the value of each HP and the resulted performance. In particular, we find that learning rate, max depth and subsample have significant impact on $\Delta\mathcal{L}$. However, the generalizability of a learner is not monopolized by a single one of them; instead, it depends on their cumulative interactions. For example, the largest $\Delta\mathcal{L}$s are observed for learners that features a large learning rate, deep base trees, combined with low subsample rate, but any of these HP settings alone does not necessarily lead to the worst case performance. In addition to this, we notice that such generalization gap is also related to dataset characteristics and weakly correlated across models, and we dissuss more about this matter in Appendix E.
Figure 5: (a) Scatter plot of $\mathcal{L}_{\text{train}}$ versus $\mathcal{L}_{\text{test}}$ for all $14,960$ configurations of XGBoost on the dataset #44059. $\bm{\lambda}^{*}_{\text{test}}$ is marked by red star , (b-e) The same plot with colormap added according to HPs values. Warmer color indicate higher values.
4.3 HPO Landscapes with Different Fidelities
Figure 2 (c) shows the low-fidelity test loss landscapes (denoted as $\mathcal{L}_{\text{test}_{\text{LF}}}$) for each model (using $10$ epochs for FCNet, and $10\%$ training data for others). From the plots, we could see that $\mathcal{L}_{\text{test}_{\text{LF}}}$ landscapes are highly consistent with $\mathcal{L}_{\text{test}}$ in terms of both structural characteristics and performance rankings. More specifically, as reflected in Figure 3, all measured FLA metrics of $\mathcal{L}_{\text{test}_{\text{LF}}}$ landscapes showed little difference compared to $\mathcal{L}_{\text{test}}$ landscapes across all studied scenarios. For performance rankings, Figure 4 (b) depicts the distribution of the $3$ similarity indicators between $\mathcal{L}_{\text{test}_{\text{LF}}}$ and $\mathcal{L}_{\text{test}}$ across all datasets for each model. We could observe a high Spearman correlation (median $>0.85$) between $\mathcal{L}_{\text{test}}$ and $\mathcal{L}_{\text{test}_{\text{LF}}}$ for all models, and the $\gamma$-set similarities between the top-$10\%$ configurations are also prominent, with medians larger than $60\%$. These imply that $R(\mathcal{L}_{\text{test}})$ and $R(\mathcal{L}_{\text{test}_{\text{LF}}})$ are highly consistent for the majority of our studied scenarios and there is large overlap between the promising regions of the two landscapes. In addition, the Shake-up scores yield low values ($\text{median}<0.1$), suggesting that on average the difference between $R(\mathcal{L}_{\text{test}})$ and $R(\mathcal{L}_{\text{test}_{\text{LF}}})$ is less than $10\%$. Additional results on FCNet, NAS-Bench-101 in Appendix D and Appendix C respectively are also consistent with our findings here.
4.4 HPO Landscapes Across Datasets
[…]
Figure 6: Importance of top-$5$ HPs as determined by functional ANOVA method for each model.
In addition to performance rankings, Figure 6 illustrates the contribution of each HP and their interactions to model performance assessed by the functional ANOVA method. The results indicate that some (combination of) HPs are typically important for many datasets for a given model. For example, learning rate consistently contributes a large portion of variance to model performance for LightGBM, and its interactions with the number of leaves and estimators are also important. These observations are similar with van Rijn & Hutter (2018), which find also conclude that certain HPs of a ML model are important for a wide spectrum of datasets by analyzing meta-data from OpenML platform.
As discussed in Section 4.1, HP loss landscapes often involve a large number of none-improvement moves, especially near the optimum. We also see that there is clear division between the promising regions and the poorly-performaning ones. Therefore, leveraging prior knowledge from previously tasks should be able to greatly expedite the searching process by means like warm-starting HPO from good configurations or carefully selecting candidate HPs and crafting the search space. More importantly, based on our results, we note that this should not be only limited to similar tasks defined under certain rules, since they may not always available. On the other hand, seemingly different tasks could still provide informative information as we see above. Our additional results for FCNet on $4$ datasets, and NAS-Bench-201 across CIFAR-10/100 as well as ImageNet datasets (Appendix C), also revealed similar highly-transferable conclusions.
5 Disuccsions and Conclusions
By conducting large-scale exploratory analysis on $1,500$ HP landscapes of $6$ ML models with over $11$M model configurations under different scenarios, this paper reveals an unified portrait of their topographies in terms of smoothness, neutrality and modality. We also show that these properties are highly transferable across datasets and fidelities, and thus provide fundamental evidence to support the effectiveness of transfer and multi-fidelity methods, which in privious practices, is mainly based on intuition. However, while our findings are observed for the majority studied scenarios, we do observe some exceptions. For example, most landscapes inspected reveal a nearly unimodal structure, but some of them can have dozens to a few hundreds of local optima with non-negligible basin sizes (e.g., FCNet). Also, there are cases when landscapes under lower fidelities or on a different task reveal very different patterns, as shown by the long tails of the similarity distributions in Figure 4. Further explorations by interrogating the relationship between dataset characteristics may provide a even more comprehensive understanding of the HPO landscape.
Our developed FLA framework in this work, has shown great potential for enabling both qualitative and quantitative understanding for a wider range of AutoML problems. While it currently relies on large-scale, pre-evaluated data points for landscape construction, we think it can be a promising direction to integrate it with existing AutoML frameworks, and thus allow for on-the-fly analysis of problem landscapes, thereby allowing it to be accessible by a broader range of stakeholders.
Appendix A Background and Related Work
Fitness Landscape of BBOPs. A central topic in optimization concerns how the intricate interplay of decision variables determine the targeted objective values. For example, geneticists and biologists are interested in the relationship between genotypes and the phenotypes and functions of organisms that may affect their evolutionary success (De Visser & Krug, 2014). In chemical physics, scientists seek to explore how the energy of a system is dependent on its various configurational or conformational states (Brooks III et al., 2001). The fitness landscape can be envisioned as a surface over the high-dimensional space as formed by decision variables, where the objective values are represented as the elevation of the landscape. The structure of this surface describes the spectrum of possible objectives values across the variable space and thus strongly influences the optimization. In addition, searching trajectories of problem solvers can be thought as strategic walks on the corresponding problem landscape. Successful applications of this metaphor to analyze black-box systems have advanced the understanding of many fields (He & Liu, 2016; Puchta et al., 2016; Shires & Pickard, 2021; Brooks III et al., 2001). Most related to our work is the fitness landscape analysis (FLA) for black-box optimization problems (BBOPs) (Malan, 2021). Many metrics that quantify the structural characteristics of fitness landscapes have been developed over the years to describe the topography of BBOP landscapes, e.g., autocorrelation, neutrality, modality, epistasis variance, evolvability, etc.
Fitness Landscape on other ML Systems. In addition to the works that focus on analyzing HP and AutoML landscapes introduced in Section 1, there are also works which applied FLA to study other ML systems, including the neural architecture search (NAS) spaces (Rodrigues et al., 2020; Traoré et al., 2021) and the HP landscapes of reinforcement learning (RL) (Eimer et al., 2023; Mohan et al., 2023). We note that these works are somewhat orthogonal to ours, as we focus on the HP landscapes of ML algorithms. In addition, there is also another line of work on investigating the loss landscapes of neural networks (Rakitianskaia et al., 2016; Rodrigues et al., 2022; van Aardt et al., 2017). These works, though seemingly similar, are focused on the neural network losses which are the objectives of weight optimizers during model training, and are thus significantly different from ours.
HPO Landscape Visualization. The ability to visualize HP loss landscapes is crucial for enabling intuitive understandings of their complex topologies, which is however, far from trivial in practice due to their high-dimensional nature. Current study methods usually bypass this issue by exmaining only two HPs each time (see, e.g., Akiba et al. (2019); Ørebæk & Geitle (2021)). However, such techqniues can not provide an unified illustration of the total interactive effects of all interested HPs. In the evolutionary computation, there are have been successful attempts on applying dimensionality reduction techniques to visualize the fitness landscape of BBOPs (Michalak, 2019; Huang & Li, 2023; Sass et al., 2022; Walter et al., 2022). Similar work has also been done in the domain of chemical physics (Shires & Pickard, 2021), where UMAP and $t$-SNE are used to visualize the so-called energy landscape.
Overfitting in Machine Learning. To empirically investigate the presence of overfitting in real-world ML applications, Roelofs et al. (2019) analyzed massive submission meta-data on the Kaggle platform, and their results suggest little evidence of overfitting due to testset reuse. However, since the Kaggle data is constituted by a diverse set of ML models with very different HP configurations that are not comparable with each other, it prohibits further analysis on which HP configurations are more likely to lead to overfitting.
Hyperparameter Importance and Interaction. Understanding which HPs influence model performance to what extend can provide valuable insights into the tuning strategy (Probst et al., 2019). To analyze importance of hyperparameters, one could either use models that are inherently interpretable, e.g., decision trees (Quinlan, 1986), or apply model-agnostic methods such as functional ANOVA (Hutter et al., 2014a; Watanabe et al., 2023b). Built on this technique, van Rijn & Hutter (2017) compared the performance of a variety of algorithms on large set of OpenML datasets and stated that the same hyperparameters were typically important across datasets.
Appendix B Details of the Landscape Analysis Framework
Figure 7: High-level conceptual workflow of our proposed HP loss landscape analysis framework.
Figure 7 illustrates the conceptual workflow of our proposed HP loss landscapes analysis framework in Section 2. From a high level, our landscape visualizations (Appendix B.1) seek to provide a general sketch of the landscape topography. The observed patterns can then be quantitatively verified via a set of dedicated FLA metrics characterizing complementary aspects of landscape properties (Appendix B.2), while landscapes visualizations, at the same time, can assist the intuitive interpretation of the numerical figures. In addition to these, we also apply $3$ dedicated similarity measures to quantify the consistency of configuration performance across landscapes (Appendix B.4).
B.1 Landscape Visualization Method
HOPE Node Embedding. To preserve the intrinsic neighborhood relationship of HP configurations, our proposed landscape visualization method first needs to learn a low-dimensional embedding for each configuration in the landscape (i.e., each node in the graph). In this paper, we choose HOPE [^1] We use the implementation in Karateclub package. node embedding method to serve this purpose, as it is able to capture asymmetric high-order proximity in directed networks. Specifically, in a directed network [^2] Most manipulations and analyses of HP loss landscapes as complex networks are grounded on NetworkX and Pandas package., if there is a directed link from vertex $v_{i}$ to vertex $v_{j}$ and from vertex $v_{j}$ to vertex $v_{k}$, it is more likely to have a link from $v_{i}$ to $v_{k}$, but not from $v_{k}$ to $v_{i}$. In order to preserve such asymmetric transitivity, HOPE learns two vertex embedding vectors $U^{s},U^{t}\in\mathbf{R}^{|V|\times d}$, which is called source and target embedding vectors, respectively. After constructing the high-order proximity matrix $S$ from $4s$ proximity measures, i.e., Katz Index, Rooted PageRank, Common Neighbors and AdamicAdar. HOPE finally learns vertex embeddings by solving the a matrix factorization problem:
$\min_{U_{s},U_{t}}\|S-{U^{s}U^{t^{T}}}\|_{F}^{2}$ (1)
UMAP Dimensionality Reduction. While HOPE (as well as any other node embedding methods) could generate vectorized embeddings for configurations, these embeddings typically are still high-dimensional and not compatible for $2$D visualization. To cope with this, we further apply UMAP [^3] We use the implementation in UMAP package. to project the HOPE embeddings into a $2$D space, which is based on three assumptions, and here we provide brief discussion on whether they hold for our case:
  • Data are uniformly distributed on Riemannian manifold. As the distribution of HOPE embeddings essentially depends on the connectivity pattern of the graph, for HP loss landscapes, we can expect this assumption to hold. This is because the number of neighbors for each node in the graph would largely stay the same based on our neighborhood definition. Therefore, there will be no significance variation of density across the graph, and the HOPE embeddings in turn should approximately have a uniform distribution.
  • The Riemannian metric is locally constant. The distances between HOPE embeddings correlate directly with the local connectivity patterns within the network. Given the stability in neighborhood structures as we discussed above, it is reasonable to presume that the local distance metrics would remain approximately constant in small regions.
  • The manifold is locally connected. HOPE embeddings can actually preserve more than just the local structure between configurations, as it is able to capture high-order proximities in the graph. We thereby expect this assumption will also hold.
Remark on Algorithm Choices. In principle, other node embedding and dimensionality reduction methods can be applied to serve our purpose. Our specific choice on HOPE is mainly because of its scalability to large-scale networks and its ability to preserve both local and global structure of the landscape. As for dimensionality reduction, Draganov et al. (2023) has made detailed theoretical comparisons between UMAP and $t$-SNE, and shows that only the normalization significantly impacts their outputs. This thens implies that a majority of the algorithmic differences can be toggled without affecting the embeddings. We choose UMAP here for its better scalability. At the end, the quality of the visualization using our method will continuously grow with the new state-of-the-art in graph represention learning and dimensionality reduction.
Remark on UMAP Hyperparameters. Since HP loss landscapes can vary a lot in terms of dimensionality and total number of configurations, there is no ‘one-size-fits-all’ setup for HPs of UMAP. However, we still provide some general guidelines for HP tuning, which mainly focus on two most important HPs, namely n_neighbors and min_dist. Specifically, n_neighbors controls the balance between local and global structure in the embedding. In general, we found that it is better to set n_neighbors to a value that is larger than the average number of neighbors for each node. For min_dist, which specifies the minimum distance between points in the low-dimensional space, we generally recommend values larger than $0.5$. The reasoning here is that we want the points to be more spread out in the low-dimensional space, and thus prevent the points to be densely distributed in local regions. However, a min_dist that is too large can also cause problems, as different parts of the landscape can get intertwined with each other.
B.2 Landscape Analysis Metrics
Algorithm 1 Best-Improvement Local Search
0:  A starting configuration $\mathbf{c}$; A neighborhood function $\mathcal{N}_{d}$; A fitness function $f$
1:  while $\mathbf{c}$ is not a local optimum do
2:     $\mathbf{c}^{\prime}_{\text{best}}=\arg\min_{\mathbf{c}^{\prime}\in\mathcal{N}_{1}(\mathbf{c})}f(\mathbf{c}^{\prime})$
3:     if $f(\mathbf{c}^{\prime}_{\text{best}})<f(\mathbf{c})$ then
4:        $\mathbf{c}\leftarrow\mathbf{c}^{\prime}_{\text{best}}$
5:     end if
6:  end while
7:  return $\mathbf{c}$
Algorithm 2 Identifying Local Optima and Their Basins
0:  The set of all configurations $\mathcal{C}$
1:  $\mathcal{V}\leftarrow\emptyset$
2:  $\mathcal{B}\leftarrow\emptyset$
3:  for all $\mathbf{c}\in\mathcal{C}$ do
4:     $\mathbf{c}^{\ell}\leftarrow\textsc{LocalSearch}(\mathbf{c})$
5:     $\mathcal{B}[\mathbf{c}^{\ell}]\leftarrow\mathcal{B}[\mathbf{c}^{\ell}]\cup\{\mathbf{c}\}$
6:  end for
7:  return $\mathcal{V},\mathcal{B}$
Assortativity coefficient. The assortativity coefficient of a network assesses the degree to which nodes tend to be connected to other nodes that are similar w.r.t. some attributes. For example, in a social network, this would mean that people tend to be friends with other people who are similar to themselves in terms of education level, income, race [^4] As one can image, the exact definition of assortativity can depend on whether the target attribute is categorical (i.e., unordered) or numerical (i.e., ordered). Here we focus on the latter as model loss is real-valued., etc. For HP loss landscape, the performance assortativity measures the extent to which HP configurations with similar performance are more likely to be neighbors to each other. Formallly, given a HP loss landscape as a directed graph, and the model loss $\mathcal{L}$ takes values $[L_{1},L_{2},\dots]$, the $\mathcal{L}$-assortativity evaluates the Pearson correlation of the measured loss between pairs of linked configurations and is measured as (Newman, 2003):
-$\mathcal{L}\text{-ast}=\frac{\sum_{i}e_{ii}-\sum_{i}a_{i}b_{i}}{1-\sum_{i}a_{i}b_{i}}$ (2)
where $e_{ij}$ is called mixing matrix entry, which represents the fraction of total edges in the network (i.e., landscape) which connects configurations having performance $\mathcal{L}(\bm{\lambda})=L_{i}$ to configurations having attribute $\mathcal{L}(\bm{\lambda})=L_{j}$. In directed networks like our case, this can be asymmetric, i.e., $e_{ij}\neq e_{ji}$. In addition, $a_{i}=\sum_{j}e_{ij}$ is the portion of edges $(\bm{\lambda}_{u},\bm{\lambda}_{v})$ such that $\mathcal{L}(\bm{\lambda}_{u})=L_{i}$ and $b_{i}=\sum_{j}e_{ji}$ is the portion of edges $(\bm{\lambda}_{v},\bm{\lambda}_{u})$ such that $\mathcal{L}(\bm{\lambda}_{v})=L_{i}$. A high $\mathcal{L}$-ast would imply that configurations with similar performance have strong tendancy to be connected and form local clusters.
Landscape Neutrality. It is often in genetics that a mutation in a single position of a DNA sequence will only lead to negligible change in the expression. Such phenomenon is known as landscape neutrality at a macro level, and for each sequence, this can be quantitatively measured by the mutational robustness (Payne & Wagner, 2019), which is the probability for such non-effective mutation to happen among all its possible mutants. Similar ideas are also applicable to HPO, where altering certain HPs may only result in subtle performance shifts. In particular, we define two neighboring configurations to be neutral if their respective performances differ less than a small fraction $\epsilon$. We choose $\epsilon=0.1\%$ in this paper, since changes below this threshold would make almost no practical meaning. We then define the neutral ratio, denoted as $\nu(\bm{\lambda})$, of a configuration $\bm{\lambda}$ as the portion of its neutral neighbors in its neighborhood. The avarage neutrality of the whole landscape is then defined as:
$\bar{\nu}=\mathbb{E}[\nu(\bm{\lambda})]=\frac{1}{|\bm{\Lambda}|}\sum_{\bm{\lambda}\in\bm{\Lambda}}\nu(\bm{\lambda})$ (3)
Neutrality Distance Correlation. While neutrality generally characterizes the expected probability for neutral moves to occur in the whole landscape, it can actually vary across regions. In particular, it is important to investigate whether it is more likely to encounter neutral moves when approaching the global optimum, as in practice, we often find a diminishing gain when tuning towards the best-possible configuration. We quantitatively assess this using the neutrality distance correlation (NDC), which measures the Pearson correlation coefficient between the neutrality of a configuration $\bm{\lambda}$ and its distance to the global optimum, $\mathrm{d}(\bm{\lambda},\bm{\lambda}^{*})$ (equation (4)). We need to note that result derived from this can be potentially affected by the choice of $\epsilon$ for neutrality. To cope with this, we also develop an alternative method for assessing NDC, which is directly based on raw loss differences between configurations. Specifically, for each adaptive walk in the landscape using best-improvement local search (Algorithm 1) that can eventually approach $\bm{\lambda}^{*}$, we measure the Pearson correlation coefficient between $\Delta\mathcal{L}$ for each pair of consecutive configurations $(\bm{\lambda}_{i},\bm{\lambda}_{i-1})$ ($i\geq 2$), and $\mathrm{d}(\bm{\lambda}_{i},\bm{\lambda}^{*})$. We calculate NDC as the average across all such walks. However, in this paper, to keep the consistency with the neutrality metric, we report the results based on the first method.
$\text{NDC}=\rho_{p}[\nu(\bm{\lambda}),\mathrm{d}(\bm{\lambda},\bm{\lambda}^{*})]$ (4)
Number of Local Optima. A configuration $\bm{\lambda}$ is said to be a local optimum if its performance is superior to any other configuration in its neighborhood, i.e., $\forall\bm{\lambda}\in{\mathcal{N}(\bm{\lambda}^{\ell})}$, we have $\mathcal{L}(\bm{\lambda}^{\ell})<{\mathcal{L}(\bm{\lambda})}$. For a unimodal landscape, there is only a global optimum configuration $\bm{\lambda}^{*}$. In constrast, multimodal landscapes have various local optima with sub-optimal performance, which can pose challenges to the optimization.
Size of Basin of Attraction. While a multimodal landscape can be difficult to optimize due to the pressence of various local optima, not all of them are equal in terms of the capability of trapping a solver. For a $2$D minimization scenario, this can be envisioned by the fact that each local optimum is located at the bottom of a ‘basin’ in the landscape surface. Configurations in each basin would eventually fall into the corresponding basin bottom (i.e., local optimum) when applying a simplest hill-climbing local search (Algorithm 1). The effort needed to escape from such a basin is direclty related to its sizes (e.g., depth and radius in a $2$D space). A local optimum with small basin size is very unlikely to cause significant obstacles to optimization, whereas the opposite is true for those featuring a dominated size of basin that is even larger than the global optimum. Formally, we define the basin of attraction $\mathcal{B}$ of a local optimum $\bm{\lambda}^{\ell}$ to be the set of all configurations from which local search converges to $\bm{\lambda}^{\ell}$, i.e., =$\mathcal{B}=\{\bm{\lambda}\in{\bm{\Lambda}}\mid\mathrm{LocalSearch}(\bm{\lambda})\rightarrow\bm{\lambda}^{\ell}\}$ (Algorithm 2). The size of $\mathcal{B}$, denoted as $s_{\mathcal{B}}$, is defined as the cardinality of the basin set as $|\mathcal{B}|$. In this study, we report the mean basin size $\bar{s}_{\mathcal{B}}$ of all local optima (except the global optimum) in the landscape.
Autocorrelation. A common metric for characterizing the smoothness of a landscape is the autocorrelation $\rho_{a}$ on a series of performance values $\mathcal{L}$. These values are extracted for configurations in a random walk $\text{RW}=\{\bm{\lambda}_{0},\bm{\lambda}_{1},\dots,\bm{\lambda}_{n}\}$ in the search space $\bm{\Lambda}$. Formally:
$\rho_{a}(k)=\frac{\mathbb{E}[(\mathcal{L}(\bm{\lambda}_{i})-\bar{\mathcal{L}})(\mathcal{L}(\bm{\lambda}_{i+k})-\bar{\mathcal{L}})]}{\mathbb{V}(\mathcal{L}(\bm{\lambda}_{i}))},\forall\bm{\lambda}_{i}\in\bm{\Lambda}$ (5)
Here, $k$ represents a lag or a step difference in the indices of configurations, and in our case we consider $k=1$ since each step in our search grids have been specifically designed to mimic the tunning strategy commonly used by human experts. For each landscape, we conduct $100$ random walks of length $100$ and average $\rho_{a}$ across all measurements to mitigate the effects of randomness.
B.3 Local Optima Network
Beyond the number of local optima and their basin sizes, a further aspect to investigate is the connectivity pattern between them. For example, an important question that we may concern is whether we can escape from a given local optimum to the global optimum?, if yes, then what is the chance of this? Local optima networks (LON) (Ochoa et al., 2008; Vérel et al., 2011), which are rooted in the study of energy landscapes in chemical physics (Stillinger, 1995), address these questions by constructing a subspace of the original fitness landscape where the nodes indicate local optima, and edges represent possible transitions between them. In particular, an improving edge can be traced from local optimum configuration $\bm{\lambda}^{\ell}_{i}$ to $\bm{\lambda}^{\ell}_{j}$ if configurations in $\mathcal{B}(\bm{\lambda}^{\ell}_{i})$ can escape to $\bm{\lambda}^{\ell}_{j}$ by applying a $2$-bit perturbation followed by local search. The edge weights $w_{i,j}$ indicate the total probability for such transitions to happen (see Algorithm 3) between two local optima. By conducting network mining on LONs (e.g., Huang & Li (2023)), we can get further insights into the distribution and connection between local optima as well as how they are potentially linked with the global optimum. However, while we believe LON could be an effective tool for analyzing complex landscapes and we incorporate this as part of our landscape analysis tool, we do not present such analyses in most scenarios in this paper since their HP loss landscapes only possess a few local optima (if not none).
Algorithm 3 Constructing Local Optima Network
0:  The set of local optima $\mathcal{V}$; The basin of attraction of each local optimum $\mathcal{B}_{\mathbf{c}^{\ell}}$; The set of all configurations $\mathcal{C}$; A neighborhood function $\mathcal{N}_{d}(\mathbf{c})$
1:  $\mathcal{E}\leftarrow\emptyset$
2:  $\mathcal{W}\leftarrow\emptyset$
3:  for all $\mathbf{c}^{\ell}\in\mathcal{V}$ do
4:     for all $\mathbf{c}^{\ell\prime}\in\mathcal{N}_{2}(\mathbf{c}^{\ell})$ do
5:        $\mathbf{c}^{\ell}_{\text{new}}\leftarrow\textsc{LocalSearch}(\mathbf{c}^{\ell\prime})$
6:        if $f(\mathbf{c}^{\ell}_{\text{new}})<f(\mathbf{c}^{\ell})$ then
7:           if edge $(\mathbf{c}^{\ell},\mathbf{c}^{\ell}_{\text{new}})$ not in $\mathcal{E}$ then
8:              $\mathcal{E}\leftarrow\mathcal{E}\cup\{(\mathbf{c}^{\ell},\mathbf{c}^{\ell}_{\text{new}})\}$
9:              $\mathcal{W}[(\mathbf{c}^{\ell},\mathbf{c}^{\ell}_{\text{new}})]\leftarrow 1$
10:           else
11:              $\mathcal{W}[(\mathbf{c}^{\ell},\mathbf{c}^{\ell}_{\text{new}})]\leftarrow\mathcal{W}[(\mathbf{c}^{\ell},\mathbf{c}^{\ell}_{\text{new}})]+1$
12:           end if
13:        end if
14:     end for
15:  end for
16:  return $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{W})$
B.4 Landscape Similarity Metrics
Spearman Correlation. It is a non-parametric measure of rank correlation which assesses how well the relationship between configuration performances in two landscapes can be described using a monotonic function. It is defined as the Pearson correlation coefficient between the performance ranks configurations in two landscapes.
Shake-up Metric. It originates from the Kaggle competition community and is designed to assess the rank changes between the public and private leaderboards of a competition (Trotman, 2019). To be specific, it quantifies the average movement of rankings from public board to the private board. For HP loss landscapes, this metric can indicate the expected rank shifts for a configuration when evaluating it on two different scenarios (e.g., change the dataset).
[…]
$\gamma$-set Similarity. It is proposed in Watanabe et al. (2023a) to assess the similarity of two tasks using the ratio of their most proninent configurations. More specifically, for two HP loss landscapes, their $\gamma$-set similarity is defined as the ratio of the intersection of the top-$\gamma$ regions to the union of them. In this paper, we consider $\gamma=10\%$.
Appendix C Analysis on NAS Loss Landscapes
C.1 NAS Benchmarks
NAS-Bench-101. It represents the first effort towards benchmarking NAS research and thus fostering reproducibility in the community. It evaluates a CNN architecture with $3$ stacked blocks, where a down-sampling is added between two consecutive blocks; A $3\times 3$ convolution is used before the main blocks, and the outputs of the main blocks are fed to an average pooling and fully connected layer. NAS-Bench-101 considers a cell-based search space which enumerates all possible configurations for each block. More specifically, the search space is formulated as a directed acylic graph (DAG) with $7$ nodes and a maximum of $9$ edges. Here, each node can represent one of the following operations: $a)$: $1\times 1$ convolution, $b)$: $3\times 3$ convolution, and $c)$: max pooling. After removing all isomorphic cells, this search space results in $423$k unique configurations. NAS-Bench-101 evaluates each of them on the CIFAR-10 dataset and records meta-data at the $\{4,12,36,108\}^{\text{th}}$ epoch.
NAS-Bench-201. It features a different skeleton compared to NAS-Bench-101, in which a residual block is applied to connect $3$ cells. Each cell here is a DAG with $4$ nodes and $6$ edges. Morever, here, operations are represented by edges, which have $5$ types: $a)$: zeroize (do nothing), $b)$: $1\times 1$ convolution, $c)$: $3\times 3$ convolution, $d)$: $3\times 3$ average pooling, $e)$: skip connection. The benchmark thus contains $5^{6}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}15,625}$ unique model architectures, with each evaluated on $3$ different datasets: $i)$: CIFAR-10, $ii)$: CIFAR-100, $iii)$: ImageNet-16-120 using $200$ epochs.
C.2 NAS Landscape Construction
Despite the cell-based search spaces of NAS benchmarks are very different from the HP ones considered in this paper, our landscape construction rountine could be easily transfered to NAS by redefining the neighborhood structure. This then demands proper encoding of the NAS configurations and the definition of a suitable distance function.
NAS-Bench-101 Neighborhood. In the original paper, a cell-encoding method based on adjacency matrices is introduced to encode configurations, which comprises two components. First, a $7\times 7$ upper-triangular binary matrix is used to indicate whether an edge exist between two nodes and thus determine the connectivity pattern of nodes. Next, the functionality of a configuration also depends on which operation is performed at each node, and this could be encoded via a vector of length $5$ (the input and output nodes are the same across archiectures and are omitted, non-existent nodes are represented by NaN). Therefore, a configuration could be specified using an adjacency matrice and a node vector. We define a neighbor of a given configuration to be the one with $1$-edit distance from it (e.g., either adding or deleting one edge, or change the operation of one node), as in the original paper. Note that not all such configurations are valid in the benchmark, as they can be isomorphic to others. The benchmark API provides built-in function to check for this.
NAS-Bench-201 Neighborhood. The encoding of configurations in this benchmark is much more straight forward. Specifically, we encode each configuration using a $6$-bit vector, where each bit specifies the operation taken at each edge, which can take $5$ categorical values. We base our neighborhood definition here on the $1$-edit distance as well, in which a $1$-bit mutant of a configuration is the one with the operation in only one edge altered.
C.3 Results on NAS Loss Landscapes
While we have conducted all analyses on both benchmarks, it would be too tedious to lay all the information here. Instead, we use NAS-Bench-101 to discuss general NAS landscape characteristics and multi-fidelity, and present the comparisons between NAS landscapes across datasets using NAS-Bench-201. Before we present our results, we also note that while the majority of configurations ($359$k out of $423$k) in NAS-Bench-101 search space come with $7$ nodes (i.e., with $5$ intermediate operations), the rest of them have less nodes. Here we mainly focus on configuragtions with $7$ nodes, since accounting all the configurations would result in many independent components in the landscape. We do not expect this to significantly affect our results.
NAS Landscape Visualization. We first visulize the NAS-Bench-101 landscape using our proposed landscape visualization method in Section 2, as shown in Figure 8. Specifically, we plot the training accuracy landscape along with test accuracy landscapes trained at $4$ different number of epochs. It is clear to see from the plot that the landscape is far from unimodal, with many local optima. However, configurations still tend to form local clusters, though the relative size of each plateau seems to be much smaller than we see for HP loss landscapes. Considering that the search space here contains nearly $30$ times more confiurations than the HP space we used in the main text, each cluster may contain thousands of configurations, inside which the landscape could still be sufficiently smooth.
Figure 8: Visualization of NAS-Bench-101 landscape using our proposed method. Here, we present the training accuracy landscapes as well as test accuracy landscapes at different recorded epochs. Color indicates rank of test accuracy, and higher values are better.
NAS Landscape Metrics. Here we report several landscape metrics for the NAS-Bench-101 landscape with respect to test accuracy at the $108$-th epoch:
  • Autocorrelation. We obtained a correlation coefficient of $0.6031$ on the landscape. This confirms our hypothesis above that the landscape is still sufficiently smooth and highly navigable, despite we observed more complex patterns in the visualizations.
  • Clustering. The accuracy-assrotativity is $0.6485$, which indicates a good level of local clustering of configurations with similar performance levels. This further confirms that the landscape is locally smooth.
  • Neutrality. Our neutrality measure, however, only yields a value of $0.075$, and suggests that most $1$-bit changes in a configuration would result in performance shift $>0.1\%$.
  • NDC. Despite the overall neutrality of the NAS landscape is low, we still observe a high NDC value of $0.7194$. This implies a strong plateau trend near the optimum, where optimizers would pay considerable more effort to gain marginal performance improvement.
  • Number of Local Optima. One of the most distinguishable property of NAS-Bench-101 is a large collection of local optima in the landscape. In fact, we found $5,908$ local optima in total (out of the $359$k configurations), which could probably make the landscape far more difficult to optimization. We will discuss more about them and their basins in the LON part.
Figure 9: Visualization of the local optima network of NAS-Bench-101 test landscapes at the $108$-th epoch, containing $5,908$ nodes. Radius of each node (i.e., local optimum) indicates size of the corresponding basin of attraction. The color indicates test accuracy, where warmer color is better. Edges indicate transition probabilities between local optima, where thicker, warmer edges imply the corresponding transition if more likely to happen. Edge directions indicate the improving direction (i.e., pointing to the fitter configuration).
NAS Landscape Across Fidelities. From Figure 8, we could see that in general the test landscapes with lower fidelities resemble the one trained with $108$ epoch. Quantitatively, the Spearman correlation between the test accuracy landscapes at $108$-th epoch and $32$-th epoch is $0.904$, with a Shake-up metric of $9.74\%$. These suggest a good general correlation between these landscapes, although there are $4$ times difference in their budget. When zooming into the $10\%$ region, the $\gamma$-set similarity is $0.64$, which is also farily good (the intesection ratio between top-$10\%$ regions is $78\%$). When we further decrease the budge by $4$ times, the Spearman correlation and Shake-up metric obtained between the $12$-th and $108$-th epoch are $0.657$ and $18.4\%$ respectively, while the $\gamma$-set similarity is $0.317$. Finally, with only $4$ epochs of training, the above metrics further change to $0.504$, $22.4\%$ and $0.164$ respectively. In general, the landscapes are still moderately correlated, but the detailed pattern can be largely distorted.
Figure 10: Visualization of NAS-Bench-201 landscape using our proposed method across datasets.
Local Optima Network Analysis.
Figure 11: Distribution of local optima test accuracy versus LON metrics.
Figure 9 shows the local optima network of NAS-Bench-101 test landscapes at the $108$-th epoch. It could be obviously seen that there are clear community (clustering) structure in the network, where local optima with large basin of attractions tend to locate at the center of each cluster, which usually feature a promising performance. To be specific, from the left plots in Figure 11, we could clearly observe that the size and radius [^5] Here, we define the radius of a basin of attraction as the expected number of local search steps needed to reach the corresponding local optimum. Since in practice, we observe that this value is often highly correlated with the size of basin of attraction, and correlate with performance in a similar manner, we only report basin size in the main text. of the basin of attraction is positively correlated with the performance of local optima (Spearman correlation $>0.55$). More importantly, as suggested by the cumulative distribution of basin size shown in Figure 12, the highly-fit local optima have a dominant size of basin of attractions in the landscape. For example, the dashed line in Figure 12 indicates that the cumulative sum of basin sizes of those local optima with $acc>94.3\%$ take $50\%$ of the total basin size (which is the number of total configurations in the landscape). Since being in the basin of a local optimum would imply that local search will eventually converge it, our result is to say that if we start local search from a random configuration in the landscape, there is $50\%$ chance that we would end up in a local optimum with $acc>94.3\%$. This is a very promising result, since we find such a level of accuracy is already better than $98.58\%$ of total configurations in the whole search space! It is also superior than $76.84\%$ of other local optima in the landscape.
Figure 12: Cumulative distribution of basin size versus local optima performance.
A natural question that follows is the efforts that we need to reach such local optima. Statistically, we find that on average, it would take $3.04$ local search steps to reach a local optimum, while the mean of the longest walk length in each basin is $6.46$ steps. This is not a huge effort, since after taking such steps, as we discussed above, we have a good chance to fall into local optimum with $acc>94.3\%$. However, we note that this is also not that trivial, since each ‘step’ here means we exhaustively search for all the neighbors of a configuration, and then select the best one to move on. Such local search technique is called a best-improvement local search. In contrast, there is also first-improvement local search, in which we take the first configuration in the neighborhood that is fitter than the current one without considering other neighbors. This in general woul require more steps to reach a local optimum, but with fewer model evaluations at each step. While local search is definitely sub-optimal compared to advanced global search strategies, finding that even such a simple technique could somehow lead to a guaranteed promising result would imply that the NAS landscape is also benign.
Beyond local search, we then continue to consider if we do fall in a local optima whose fitness is not that satisfactory, then what is the chance that we can somehow escape from it?. The two plots at the right panel of Figure 11 show the distribution of escape rate and improve rate correlate with test accuracy. Here, the escape rate is the chance that, a configuration in the basin of a local optimum $\mathbf{c}^{\ell}$, after applying a $2$-bit perturbation, would converge to a different local optimum $\mathbf{c}^{\ell}_{\text{new}}$. On the other hand, the improve rate further restricts that the new local optimum should have better performance compared to the current one. From Figure 11, we could clearly see observe the majority of the local optima feature a escape rate of larger than $50\%$, which suggests that most local optima are not that difficult to escape from. For improve rate, we observe a good correlation with test accuracy, where it is more easy to find an improving move for a poorly-performed local optimum. Unfortunately, for those that already have promising performance, there is only little chance to transit to a better basin using $2$-bit perturbation.
C.4 Results on NAS-Bench-201
We visualize the loss landscapes of NAS-Bench-201 on the $3$ datasets, namely CIFAR-10/100 and ImageNet, in Figure 10. In general, we see that results on these $3$ tasks reveal strong consistency to each other (Spearman correlation $>0.95$), which conforms with our findings on HP loss landscapes.
Appendix D Details About FCNet HP Loss Landscapes
Figure 13: Local optima network for the $100$ local optima in the FCNet landscape on parkinsons telemonitoring dataset. The left plot shows the full-view of the network, where node size represents the size of the corresponding basin of attraction. Node color indicates the performance of the model, which is labeled as rank values for each node, and the global optimum has a rank of $1$. The right plot shows the neighborhoods of the global optimum, while other nodes are omitted.
Table 3: Landscape Features of FCNet on Different Datasets
 Dataset Autocorr. Assortativity Neutrality NDC #LO Mean Basin Size
Parkinsons Tele. $0.4683$ $0.6232$ $0.4317$ $0.6368$ $100$ $589$
Protein Struc. $0.4896$ $0.5789$ $0.4746$ $0.7411$ $128$ $385$
Naval Prop. $0.5720$ $0.5298$ $0.4091$ $0.6433$ $347$ $159$
Slice Local. $0.4887$ $0.6134$ $0.4427$ $0.8216$ $24$ $2372$
D.1 Landscape Features
Figure 14: Spearman correlation between LON features ($i)$: mean basin size, $ii)$: mean basin radius, $iii)$: maximum basin radius, $iv)$: in-coming strengths, $v)$: out-going strengths, $vi)$: closeness centrality, $vii)$: eigenvector centrality, $viii)$: pagerank centrality) and local optima performance for FCNet across datasets.
From the visualization of FCNet landscapes shown in Figure 2, we can conclude that they generally follow the same characteristics as we discussed in Section 4.1. Most of such observations are supported by FLA metrics reported in Table 3. However, we note that all four landscapes seem to reveal multimodal structures, as there are dozens to hundreds of local optima with relative large basins (up to a mean size of $2,372$). Similar observations have been made by Pushak & Hoos, who speculated that the reasons could be FCNet scenarios fall into the over-parameterized regime, which is different from the other scenarios. While we also agree with this reasoning, we seek to conduct further analysis on the local optima using the local optima network (LON) (Appendix B.3). Figure 13 (left) shows the LON of FCNet on the parkinsons telemonitoring dataset. It can be seen that in general, local optima with better performance tend to feature a large basin of attraction and high in-coming strenghts (i.e., in LON context, the total probability for other local optima to reach it). This implies that, despite the pressence of many other local optima, the global optima still plays a pivotal role in the connectivity of the network (see Figure 13, right), which is not the worst case that one may encounter (e.g., a global optimum with tiny basin locates at a secluded region that the optimizer can hardly find.)
Appendix E Landscape Similarity & Datasets Characteristics
Figure 15: Spearman correlation between landscape similarity and dataset characteristics. Left: similarity between training and test loss landscapes for different models versus dataset characteristics. Right: similarity between landscape on different dataset for different models versus ratio of dataset characteristics. Here, dataset size is defined as the product of the number of instances and features.
While in general, we find that HP loss landscapes studied in this paper share various common characteristics, the detailed topography can vary with dataset that the model is trained on. How characteristics of datasets could affect the landscape is an interesting problem to explore in more detail. In addition, we also hypothesis that the generalization gap of a model can also depend on the dataset, in addition to its HP setting. To investigate these, we conduct additional experiments on the $57$ tabular datasets to analyze the relationship between landscape similarity and dataset characteristics.
We first find in the left plot of Figure 15 that the similarity between test and training loss landscapes is positively correlated across models. It implies that on certain datasets, all the $4$ models are more prone to overfit, whereas the opposite could be true for other datasets. This then verifies our hypothesis that overfitting not only depends on the model HPs, but also the dataset itself. To further explore which properties of the dataset could potentially contribute to this, we investigated the correlation between train-test landscape similarity and the number of instances & features of each dataset, and their product. The results (also in Figure 15 (left)) indicate that the degree of overfitting is correlated with all $3$ dataset size measures. In particular, for most models, it would be more likely to encounter overfitting on larger datasets.
We then proceed to investigate whether correlations between HP loss landscapes induced on different datasets are related to the relative size of the datasets. From Figure 15 (right), it is clear to see that pairwise landscape similarities are obviously correlated across models, implying that all the models are likely to induce very different (or the opposite) landscapes on certain pairs of datasets. We could also see that these pairwise similarities are again correlated with the differences in dataset sizes. In particular, for datasets that have very different sizes, the resulting HP loss landscapes also tend to be different from each other.
However, we note that the correlations reported in both scenarios are somewhat weak, and the reasons could be two folds. First, the choice of datasets is only a partial factor contributing to landscape similarity, many other factors like model HPs, training settings can also important roles here. Second, the dataset meta-features we used here are rather naive, and there could be more comprehensive features for characterizing dataset properties, e.g., Feurer et al. (2015b) had used a collection of $46$ features to assess dataset similarity. However, this is beyond the scope of this work and we leave it to future works.
Appendix F Details of the Experimental Setup
F.1 Search Spaces
In this subsection, we elaborate the principles that we follow in designing the search spaces for each model. We also provide the detailed hyperparameter grid space for them in Table 4 (XGBoost), Table 5 (DT), Table 6 (RF), Table 7 (RF), Table 8 (CNN) and Table 9 (FCNet). The top-level principle we follow is to include commonly used HPs and exclude unimportant ones, while keeping a good balance between search space coverage and computational cost. We first determine the list of HPs to be considered, and to this end:
  • We surveyed commonly used HPs in the practice (e.g., van Rijn & Hutter identified important HPs of several models using large-scale meta-data from OpenML) and HPO literature (e.g., the search spaces used in HPOBench (Eggensperger et al., 2021)). For CNN, we also refer to the design of NAS search spaces (Chitty-Venkata et al., 2023).
  • We also run preliminary experiments by fixing all except but one HP to their default values, and vary a single HP using a wide range of values. This allows us to estimate the influence of each HP on model performance without conducting large-scale search.
We then combine the knowledge obtained in these precedures to design a rough HP list to be studied for each model. We then proceed to determine the domain and granularity of each HP, for which we bear the following considerations:
  • The domain of each HP should be large enough to cover the range of values that are commonly used in practice, while keep a balance with the computational cost. For example, we can not afford to search for XGBoost with thousands of base learners.
  • Given the limited computational budget, there should be more number of bins for HPs that are more important, e.g., learning rates.
  • Given the limited computational budget, we would prefer slightly removing $1$ or $2$ less important HPs, than having a search grid with many real-valued HPs have only $2$ to $3$ bins.
Following these principles, our final search spaces generally contain $5$ to $8$ HPs, with total number of configurations ranging from $6,480$ to $24,200$. We consider these search spaces representative of the real-world practice and thus form a good basis for landscape analysis.
Table 4: XGBoost HP Grid Space ($14,960$ Configurations)
 Hyperparameter Grid Values Count
learning_rate $[1\mathrm{e}^{-3},1\mathrm{e}^{-2},3\mathrm{e}^{-2},5\mathrm{e}^{-2},7\mathrm{e}^{-2},1\mathrm{e}^{-1},\dots 5\mathrm{e}^{-1}]$ $11$
subsample $[0.2,0.4,0.6,0.8,1]$ $5$
max_depth $[4,5,6,\dots,20]$ $17$
max_bin $[256,512,1024,2048]$ $4$
n_estimators $[100,200,300,500]$ $4$
Table 5: DT HP Grid Space ($24,200$ Configurations)
 Hyperparameter Grid Values Count
splitter ["best", "random"] $2$
min_samples_split $[2,4,\dots,20]$ $10$
max_samples_leaf $[1,2,4,\dots,20]$ $11$
max_depth $[5,10,\dots,50,\texttt{None}]$ $11$
max_features $[0.1,0.2,\dots,1.0]$ $10$
Table 6: RF HP Grid Space ($11,250$ Configurations)
 Hyperparameter Grid Values Count
min_samples_split $[2,5,10,15,20]$ $5$
max_samples_leaf $[1,5,10,15,20]$ $5$
Bootstrap [True, False] $2$
max_features $[0.2,0.4,\dots,1.0]$ $5$
n_estimators $[50,100,150,200,300]$ $5$
max_depth $[10,15,\dots,50,\texttt{None}]$ $9$
Table 7: LightGBM HP Grid Space ($13,440$ Configurations)
 Hyperparameter Grid Values Count
learning_rate $[1\mathrm{e}^{-2},5\mathrm{e}^{-2},7\mathrm{e}^{-2},1\mathrm{e}^{-1},\dots 5\mathrm{e}^{-1},]$ $10$
boosting_type ["gbdt", "dart"] $2$
max_depth $[5,\dots,30,\texttt{None}]$ $16$
n_estimators $[50,75,100,150,200,250,300]$ $7$
num_leaves $[10,20,30,40,50,100]$ $6$
Table 8: CNN HP Grid Space ($6,480$ Configurations)
 Hyperparameter Grid Values Count
act_fn ["gelu", "relu", "tanh"] $3$
batch_norm [True, False] $2$
learning_rate $[1\mathrm{e}^{-4},5\mathrm{e}^{-4},1\mathrm{e}^{-3},3\mathrm{e}^{-3},5\mathrm{e}^{-3}]$ $5$
batch_size $[128,256]$ $10$
drop_out $[0.0,0.1,0.25,0.5]$ $4$
width_linear $[256,512,1024]$ $3$
width_conv $[32,64,128]$ $3$
#conv_block $[2,4,6]$ $3$
Table 9: FCNet HP Grid Space ($62,208$ Configurations)
 Hyperparameter Grid Values Count
initial_LR $\{5\mathrm{e}^{-4},1\mathrm{e}^{-3},5\mathrm{e}^{-3},1\mathrm{e}^{-2},1\mathrm{e}^{-2},.1\mathrm{e}^{-1}\}$ $6$
batch_size $\{8,16,32,64\}$ $4$
LR_schedule {cosine, fix} $2$
activation_layer_1 {relu, tanh} $2$
activation_layer_2 {relu, tanh} $2$
layer_1_Size $\{16,32,64,128,256,512\}$ $6$
layer_2_Size $\{16,32,64,128,256,512\}$ $6$
dropout_Layer_1 $\{0.0,0.3,0.6\}$ $3$
dropout_Layer_2 $\{0.0,0.3,0.6\}$ $3$
F.2 Datasets
Here we provide basic information regarding the $5$ groups of datasets used for this study, including: $i)$: numerical regression (Table 10), $ii)$: numerical classification (Table 11), $iii)$: categorical regression (Table 12), $iv)$: categorical classification (Table 11), and $v)$: image classifcation (Table 14).
Table 10: Numerical regression
 OpenML ID Dataset Name #Samples #Features
44132 cpu_act $8,192$ $21$
44133 pol $15,000$ $26$
44134 elevators $16,599$ $16$
44136 wine_quality $6,497$ $11$
44137 Ailerons $13,750$ $33$
45032 yprop_4_1 $8,885$ $42$
44138 houses $20,640$ $8$
44139 house_16H $22,784$ $16$
45034 delays_zurich_transport $5,465,575$ $9$
44140 diamonds $5,3940$ $6$
44141 Brazilian_houses $10,692$ $8$
44142 Bike_Sharing_Demand $17,379$ $6$
44143 nyc-taxi-green-dec-2016 $581,835$ $9$
44144 house_sales $21,613$ $15$
44145 sulfur $10,081$ $6$
44146 medical_charges $163,065$ $5$
44147 MiamiHousing2016 $13,932$ $14$
44148 superconduct $21,263$ $79$
Table 11: Numerical classification
 OpenML ID Dataset Name #Samples #Features
44120 electricity $38,474$ $7$
44121 covertype $566,602$ $10$
44122 pol $10,082$ $26$
44123 house_16H $13,488$ $16$
44125 MagicTelescope $13,376$ $10$
44126 bank-marketing $10,578$ $7$
45019 Bioresponse $,43$ 434 $419$
44128 MiniBooNE $72,998$ $50$
45020 default-of-credit-card-clients $13,272$ $20$
44129 Higgs $940,160$ $24$
44130 eye_movements $7,608$ $20$
45022 Diabetes130US $71,090$ $7$
45021 jannis $57,580$ $54$
45089 credit $16,714$ $10$
45028 california $20,634$ $8$
Table 12: Categorical regression
 OpenML ID Dataset Name #Samples #Features
45041 topo_2_1 $8,885$ $255$
44055 analcatdata_supreme $4,052$ $7$
44056 visualizing_soil $8,641$ $4$
45045 delays_zurich_transport $5,465,575$ $12$
44059 diamonds $53,940$ $9$
45046 Allstate_Claims_Severity $188,318$ $124$
44061 Mercedes_Benz_Greener_Manufacturing $4,209$ $359$
44062 Brazilian_houses $10,692$ $11$
44063 Bike_Sharing_Demand $17,379$ $11$
45047 Airlines_DepDelay_1M $1,000,000$ $5$
44065 nyc-taxi-green-dec-2016 $581,835$ $16$
45042 abalone $4,177$ $8$
44066 house_sales $21,613$ $17$
45043 seattlecrime6 $52,031$ $4$
45048 medical_charges $163,065$ $5$
44068 particulate-matter-ukair-2017 $394,299$ $6$
44069 SGEMM_GPU_kernel_performance $241,600$ $9$
Table 13: Categorical classification
 OpenML ID Dataset Name #Samples #Features
44156 electricity $38,474$ $8$
44157 eye_movements $7,608$ $23$
44159 covertype $423,680$ $54$
45035 albert $58,252$ $31$
45039 compas-two-years $4,966$ $11$
45036 default-of-credit-card-clients $13,272$ $21$
45038 road-safety $111,762$ $32$
Table 14: Image Classification
 Dataset #Train Samples # Test Samples #Categories Input Size
MNIST $60,000$ $10,000$ $10$ $28\times 28$
Fashion-MNIST $60,000$ $10,000$ $10$ $28\times 28$
K-MNIST $60,000$ $10,000$ $10$ $28\times 28$
Q-MNIST $60,000$ $60,000$ $10$ $28\times 28$
CIFAR-10 $50,000$ 10,$00,0$ 000 $10$ $32\times 32$
CIFAR-100 $50,000$ $10,000$ $10$ $32\times 32$

### v2

@@
-On the Hyperparameter Loss Landscapes of Machine Learning Algorithms
+On the Hyperparameter Loss Landscapes of Machine Learning Models: An Exploratory Study
 Abstract
-Despite the recent success in a plethora of hyperparameter optimization (HPO) methods for machine learning (ML) models, the intricate interplay between model hyperparameters (HPs) and predictive losses (a.k.a fitness), which is a key prerequisite for understanding HPO, remain notably underexplored in our community.
-This results in limited explainability in the HPO process, rendering a lack of human trust and difficulties in pinpointing algorithm bottlenecks.
-In this paper, we aim to shed light on this black box by conducting large-scale fitness landscape analysis (FLA) on $1,500$ HP loss landscapes of $6$ ML models with more than $11$M model configurations, across $67$ datasets and different levels of fidelities.
-We reveal the first unified, comprehensive portrait of their topographies in terms of smoothness, neutrality and modality.
-We also show that such properties are highly transferable across datasets and fidelities, providing fundamental evidence for the success of multi-fidelity and transfer learning methods.
-These findings are made possible by developing a dedicated FLA framework that incorporates a combination of visual and quantitative measures.
-We further demonstrate the potential of this framework by analyzing the NAS-Bench-101 landscape, and we believe it is able to faciliate fundamental understanding of a broader range of AutoML tasks.
+Previous efforts on hyperparameter optimization (HPO) of machine learning (ML) models predominately focus on algorithmic advances, yet little is known about the topography of the underlying hyperparameter (HP) loss landscape, which plays a fundamental role in governing the search process of HPO.
+While several works have conducted fitness landscape analysis (FLA) on various ML systems, they are limited to properties of isolated landscape without interrogating the potential structural similarities among them.
+The exploration of such similarities can provide a novel perspective for understanding the mechanism behind modern HPO methods, but has been missing, possibly due to the expensive cost of large-scale landscape construction, and the lack of effective analysis methods.
+In this paper, we mapped $1,500$ HP loss landscapes of $6$ representative ML models on $63$ datasets across different fidelity levels, with $11$M$+$ configurations.
+By conducting exploratory analysis on these landscapes with fine-grained visualizations and dedicated FLA metrics, we observed a similar landscape topography across a wide range of models, datasets, and fidelities, and shed light on several central topics in HPO.
 1 Introduction
-In the past decade, considerable efforts have been invested in developing hyperparameter optimization (HPO) techniques to automate the laborious task of hyperparameter (HP) tunning for machine learning (ML) models.
-Many successful approaches (Bergstra et al., 2011; Snoek et al., 2012; Hutter et al., 2011; Srinivas et al., 2010; Karnin et al., 2013; Li et al., 2017; Falkner et al., 2018; Awad et al., 2021) have significantly advanced this field, and they have been empirically shown to outperform both manual configuration (Hutter et al., 2019; Bischl et al., 2023; Santu et al., 2022; Yang & Shami, 2020) and random search (Bergstra & Bengio, 2012).
-HPO is often casted as a black-box optimization problem (BBOP), where the goal is to search for an HP configuration $\bm{\lambda}\in\bm{\Lambda}=\Lambda_{1}\times\dots\Lambda_{n}$ with an objective value $\mathcal{L}(\bm{\lambda})$ as small as possible without any explicit knowledge of the ML loss function $\mathcal{L}:\bm{\Lambda}\to\mathbb{R}$.
-Existing methods (see examples above) to this end essentailly comprises $3$ key components: $i)$: a search space, $ii)$: an optimization strategy, and $iii)$: model evaluation.
-While the development of both efficient searching mechanisms and evaluation strategies has received considerable attention in recent years, the intricate interplay between model HPs and predictive losses, which plays a pivotal role in understanding HPO problems, remain notably underexplored.
-Such lack of knowledge in turn hampers the transparency and explainability (Dwivedi et al., 2023) of HPO solvers, which often function as black-boxes as well.
-Consequently, this results in limited human trust in HPO methods and hinders their wide-spread application (Drozdal et al., 2020; Bischl et al., 2023) Unfortunately, given the high-dimensional, hybrid nature of HP configuration space, it is far from trivial to open up such black box.
-The fitness landscape metaphor, which was pioneered by Wright in 1932 in evolutionary biology, has been widely recognized as a powerful tool for analyzing BBOPs in the evolutionary computation community (Malan, 2021).
-It can be envisioned as a (hyper-)surface as formed by objective values, over the high-dimensional space of possible configurations (Romero et al., 2013).
-Since the 90s, a plethora of fitness landscape analysis (FLA) methods have been developed to conduct exploratory analysis on landscape characteristics of BBOPs (Zou et al., 2022).
-Such methods are useful, as they are able to extract landscape measures that are indicative of problem difficulty and how a certain search mechanism would perform on it (Smith-Miles & Lopes, 2012; Hutter et al., 2014b; Qasem & Prügel-Bennett, 2010).
-This knowledge would then advance the understanding of the problem characteristics (Huang & Li, 2023), assist the selection and configuration of problem solvers (Kerschke et al., 2019; Schede et al., 2022), navigate the design of new algorithms (Qasem & Prügel-Bennett, 2010), and enhance the explainability and trust for optimization (Thomson et al., 2023).
-Figure 1: $2$D visualization of HP loss landscapes for: (a-d) CNN on $9$ HPs and $6,480$ configurations, (e-f) XGBoost regressor on $5$ HPs and $14,960$ configurations, under different scenarios.
-Colors indicate ranks of configurations (lower values are better).
-Coordinates are projected using UMAP.
-Recently, the use of FLA in analyzing HP and the related AutoML loss landscapes has also received considerable attention.
-Various works have studied diverse structural characteristics of these landscapes including neutrality, modality, and fitness distance correlation (e.g., Pushak & Hoos (2022); Teixeira & Pappa (2022); Pimenta et al.
-(2020); Schneider et al.
-(2022)).
-However, such works suffer from limited setups and fail to interrogate the connection between landscape characteristics and the success of HP optimizers, which often run in a wide range of scenarios (e.g., different models, datasets and fidelities).
-It remains unclear whether the HP loss landscapes induced on different settings share certain characteristics or patterns, and how such commonalities could be potentially exploited by HP optimizers.
-On the other hand, we argue that existing analytical methods are insufficient to provide a comprehensive analysis and comparision on HP loss landscapes since:
-  • The ability to visualize landscapes is crucial for enabling intuitive understandings of their complex topologies (Michalak, 2019).
-However, HP loss landscapes are notoriously difficult to visualize in a human-comprehensible fashion due to their high-dimensional nature.
-Some existing methods address this problem by plotting only one or two HPs each time (e.g., Friedman (2001); Akiba et al.
-(2019)), which fail to provide an integrated view of the global landscape structure.
-Other works applied dimensionality reduction techniques to project the landscape into $2$D space (e.g., Michalak (2019); Biedenkapp et al.
-(2018); Walter et al.
-(2022)), but the resulting plot is not able to preserve the overall topography as well as neighborhood structure of the landscape.
-  • There is no tangible method for quantifying the similarity between different HP loss landscapes.
-Despite general FLA metrics could already capture informative landscape characteristics, practices in automated algorithm selection demonstrate that domain-specific metrics are also crucial as a complementary source of information for better characterizing the target problem (Smith-Miles, 2008; Smith-Miles & Lopes, 2012).
-However, none of the prior works have considered such perspectives when comparing HP loss landscapes.
-The overarching goal of this paper is to gain an integral view of the HP loss landscapes induced on different scenarios and thereby provide new insights to the community.
-To this end, we develop a dedicated landscape analysis framework to enable comprehensive analysis and comparisions among HP loss landscapes.
-It incorporates 1: a novel neighborhood-aware HP loss landscape visualization method applicable to high-dimensions, 2: a series of FLA metrics quantifying landscape structural characteristics, and 3: $3$ similarity metrics that leverage rankings of HP configurations to allow for informative landscape similarity quantification in the HPO context.
-Through empirical analysis on $1,500$ landscapes across $6$ ML models and $67$ datasets with more than $11$ million configurations, we are ambitious to advance the understanding of the following four fundamental HPO scenarios:
-HP Landscapes of Test Loss Versus Training Loss.
-‘Overfitting’ is one of the biggest interests and concerns in the ML community (Ng, 1997; Caruana et al., 2000; Recht et al., 2019; Belkin et al., 2018; Roelofs et al., 2019; Ishida et al., 2020).
-However, there is a lack of in-depth understanding on how test loss correlates with training loss across a broad HP landscape, and what specific properties distinguish regions that generalize well from poorly generalized ones.
-In this paper, by using our fitness landscape analysis framework, we find that the test loss landscapes resemble their training counterparts in terms of both structural characteristics and performance rankings (see, e.g., Figure 1 (a) versus (b)), and configurations with small training error are likely to achieve a mild generalization error.
-However, significant discrepancies can also occur (see, e.g., Figure 1 (e) versus (f)) depending on both the choice of certain HP combinations and the dataset at hand.
-In such cases, struggling to reduce the training loss has little or even negative effect to refining the generalization loss.
-HP Loss Landscapes Across Fidelities.
-Given the time-demanding nature of model evaluation, multi-fidelity HPO methods (Karnin et al., 2013; Kandasamy et al., 2016; Li et al., 2017; Kandasamy et al., 2017; Falkner et al., 2018; Awad et al., 2021) have achieved prominent performance by more efficient resource allocation.
-However, the validity of their underpinned assumption, i.e., the ranking of configuration performance would stay close to the ground truth under fidelities (Bischl et al., 2023), remains unclear (Pushak & Hoos, 2022).
-Our empirical results are highly inspiring to support such assumption and show that landscapes with lower fidelities are highly consistent with full-fidelity landscapes w.r.t.
-both structural characteristics and performance ranks (Figure 1 (c)).
+Hyperparameter optimization.
+In the past decade, considerable efforts have been invested in developing hyperparameter optimization (HPO) techniques to automate the laborious task of hyperparameter (HP) tunning of machine learning (ML) models, and many successful methods (e.g, [1, 2, 3, 4, 5, 6, 7, 8, 9], and see surveys [10, 11, 12, 13]) have significantly advanced this field.
+HPO is often cast as a black-box optimization problem, where the goal is to efficiently search for an HP configuration $\bm{\lambda}\in\bm{\Lambda}=\Lambda_{1}\times\dots\times\Lambda_{n}$ for an ML algorithm $\mathcal{A}$ that yields a loss value $\mathcal{L}(\bm{\lambda})$ as small as possible on a given dataset $\mathcal{D}$.
+By “black-box” it means that while we can evaluate $\mathcal{L}(\bm{\lambda})$ for any $\bm{\lambda}\in\bm{\Lambda}$, we have no access to the explicit form of the loss function $\mathcal{L}:\bm{\Lambda}\to\mathbb{R}$, nor other information like the gradients or the Hessian.
+The lack of problem understanding.
+This lack of knowledge on $\mathcal{L}$ in turn renders HPO solvers to function as black boxes as well, and many of them are often based on rough intuition and assumptions.
+Albeit their prominent empirical output performance, there is a lack of understanding of the inside mechanism to support their reliability and validity.
+Consequently, this poses challenges for pinpointing bottlenecks of the optimizers and thus further improvement.
+Even worse, such lack of transparency also hinders the development of human trust in HPO, hampering its wide-spread application [11, 14].
+Addressing these requires a deep understanding of the global topography of the loss function surface, which is, however, non-trivial due to its black-box nature.
+Also, as pointed out by a position paper [15] at ICML’24 [$\cdots$ in major ML venues, there is little incentive to engage in exploratory ML research].
+Fitness landscape.
+The fitness landscape metaphor, which was pioneered by Sewall Wright nearly a century ago [16], is a fundamental concept in evolutionary biology [17, 18, 19, 20] and has been widely adapted for understanding black-box systems across various domains including combinatorial optimization [21, 22, 23] and physical chemistry [24, 25].
+Intuitively, it can be conceived as a (hyper-)surface as formed by fitness ($\mathcal{L}$ in our context) values across a high-dimensional configuration space (Fig.
+1).
+Each spatial location in this landscape represents a configuration (i.e., $\bm{\lambda}\in\bm{\Lambda}$), with its elevation indicating the fitness (i.e., $\mathcal{L}(\bm{\lambda})$) and optimal configurations sitting on the valleys.
+As an optimization process is akin to navigating downhill towards these valleys (e.g., uphill trajectories in Fig.
+1), the topography of the underlying landscape contains essentially all the information needed for understanding system properties.
+Over the past several decades, a plethora of fitness landscape analysis (FLA) techniques [19, 26, 27] have been developed to characterize this topography, e.g., ruggedness [28], modality [29], neutrality [30], among many others.
+Figure 1: Fitness landscape of an illustrative $2$D problem.
+Current gaps.
+While several works [31, 32, 33, 34, 35] have explored various properties of the HP loss landscapes of ML models, they are often in a restricted scope (e.g., only a handful of landscapes), which limits the generalizability of their findings.
+More importantly, prior works often focus on individual landscapes and fail to interrogate the potential structural similarities [21] among different landscapes induced on distinct models, datasets, training fidelities, and loss functions, which are concerned by many central topics in HPO, (e.g., transfer learning [36] and multi-fidelity optimization [37]).
+Yet, the exploration of such similarities does not simply equate to an additive extension of prior works, but further requires both qualitative and quantitative FLA methods that could enable comprehensive comparisons among different HP loss landscapes.
+Unfortunately, there are no bespoke FLA metrics for this purpose, and existing landscape visualization approaches [38, 39, 40, 41, 42] struggle to tackle high-dimensionality while preserving neighborhood structure at the same time.
+[…]
+  • HP loss landscapes on training versus test data.
+Overfitting is one of the biggest concerns in ML [43, 44, 45, 46, 47, 48].
+While there has been empirical investigation on adaptive overffiting [45] with scattered data points from Kaggle competitions [47], there is a lack of more structured understanding, e.g., how the topography of the training loss landscape resemble that of its test counterpart?
+  • HP loss landscapes with different fidelities.
+Given the time-demanding nature of model evaluation, multi-fidelity HPO methods [6, 49, 7, 50, 8, 9] have achieved prominent performance by more efficient resource allocation.
+However, while there exists an implicit assumption: the topography of HP loss landscapes with lower fidelities would stay close to the groundtruth [31, 11], it has only been indirectly verified by algorithmic performances, without landscape-level evidence.
   • HP loss landscapes across datasets.
-Leveraging priors obtained from previous tasks to expedite the learning process for new tasks is another promising direction for HPO (Feurer et al., 2015b; Bardenet et al., 2013; Wistuba et al., 2015b; Kim et al., 2017; Rakotoarison et al., 2022; Wistuba et al., 2015c; Vanschoren, 2018; Swersky et al., 2013).
-These approaches are grounded in the hypothesis that ‘knowledge’ about HPO landscapes—such as configuration quality, hyperparameter importance and their interactions (Hutter et al., 2014a; Watanabe et al., 2023b; Probst et al., 2019; van Rijn & Hutter, 2017)—is correlated across related datasets when defined under a certain distance measure.
-A natural question that arises is whether this knowledge remains consistent when applied to a broader range of tasks.
-Our results on a diverse set of $67$ datasets show that performance rankings, HP importance and their interactions, are largely shared across tasks (Figure 1 (d)).
+Leveraging priors obtained from previous tasks to expedite the learning process for new tasks is another promising direction for HPO [36, 51, 52, 53, 54, 55, 56, 57].
+However, despite the prevalence of such transfer learning methods, there is a lack of understanding of how HP loss landscapes of the same model induced on different datasets would compare to each other?
   • HP loss landscapes across models.
-Methods rooted in Bayesian optimization (e.g., Snoek et al.
-(2012); Bergstra et al.
-(2011); Feurer et al.
-(2015a); Kandasamy et al.
-(2017)) and search space pruning (e.g., Wistuba et al.
-(2015a); Perrone & Shen (2019); Wang et al.
-(2020)) implicitly assume that the quality of configurations is locally correlated rather than a full randomness.
-This is also true for meta-heuristic methods (e.g., Friedrichs & Igel (2005); Lessmann et al.
-(2005); Cawley (2001); Guo et al.
-(2008)), which are even more dependent on specific landscape properties.
-While it may seem intuitive that HP loss landscapes would differ depending on the target ML model, in practice the fact is often that common HPO methods perform robustly for different models.
-This implies that, despite superficial differences, the general family of HP loss landscapes may share certain inherent patterns/properties.
-We verified this hypothesis by synthesizing the results from diverse FLA metrics characterizing HP loss landscape geometry combined with visual inspections (see, e.g., Figure 1 (a, e)).
-The results gathered from $1,500$ landscapes of $6$ ML models under different scenarios, reveal a universal picture of the HP loss landscapes.
-In this picture, HP loss landscapes are smooth, nearly unimodal, containing a large number of neutral regions; configurations with similar performance are locally clustered; the landscape becomes flatter around the optimum configuration.
-2 HPO Landscape Analysis Methods
-This section introduces our analytical framework developed to enable exploratory analysis on different HP loss landscapes and perform comparisons between them.
-Due to the page limit, we only provide a brief sketch of these methods here while more detailed discussions are in Appendix B.
-HP Loss Landscape Construction.
-The HP loss landscape can be formulated as a triplet $\langle\bm{\Lambda},\mathcal{L},\mathcal{N}\rangle$ with three ingredients: $i)$ a search space $\bm{\Lambda}$ of feasible configurations that consists of pre-evaluated, discretized grid values (see Appendix F.1), $ii)$ a ML loss function $\mathcal{L}:\bm{\lambda}\to\mathbb{R}$, and $iii)$ a neighborhood structure $\mathcal{N}$ that specifies which configurations are neighbors to each other.
-Note that the form of $\mathcal{N}$ depends on a distance function $\mathrm{d}:\bm{\lambda}\times\bm{\lambda}\to\mathbb{N}$.
-Following Pushak & Hoos (2022), we define all categorical values of a HP to be distance $1$ from each other (i.e., the values are non-ordinal).
-For a numerical HP, we define the distance between two values to be the number of steps between them on the grid used for discretization.
-Such distance measure is able to mimic the tunning strategy of human experts when combined with elaborately designed grid values.
-Based on this, the total distance between two configurations $\bm{\lambda}_{i}$ and $\bm{\lambda}_{j}$ is then sum of the distances between the respective pairs of HP values, and we say they are neighbors to each other (i.e., $\bm{\lambda}_{j}\in\mathcal{N}(\mathcal{\bm{\lambda}}_{i})$), if $\mathrm{d}(\bm{\lambda}_{j},\bm{\lambda}_{i})=1$.
-Finally, the HPO landscape is constructed as a directed graph where the vertices are HP configurations and an improving edge $e_{i,j}\in\mathcal{E}$ is traced from $\bm{\lambda}_{i}$ to $\bm{\lambda}_{j}$ if $\bm{\lambda}_{j}\in\mathcal{N}(\bm{\lambda}_{i})$ and $\mathcal{L}(\bm{\lambda}_{j})<\mathcal{L}(\bm{\lambda}_{i})$.
-We say that a configuration $\bm{\lambda}_{\ell}$ is a local optimum if $\forall\bm{\lambda}^{\prime}\in\mathcal{N}(\bm{\lambda}_{\ell})$, we have $\mathcal{L}(\bm{\lambda}_{\ell})<\bm{\lambda}^{\prime}$.
-In addition, we say that $\bm{\lambda}_{j}$ is a neutral neighbor of $\bm{\lambda}_{i}$ if their performance difference is negligible ($\leq 1$‰).
-Landscape Visualization.
-We develop a first-of-its-kind, highly interpretable method for visualizing the topography of high-dimensional HP loss landscapes by leveraging graph representation learning (Hamilton, 2020) combined with dimensionality reduction (Draganov et al., 2023) techniques.
-Specifically, we first extracted low-dimensional features for each node in the graph.
-To this end, we use HOPE (Ou et al., 2016) node embedding method because it could preserve high-order proximities between configurations.
-Then, we compress the obtained feature vectors into $2$ components using the UMAP (McInnes & Healy, 2018) algorithm, and thus allowing configurations to be laid out in $2$D scatter plot.
-To further refine the interpretability of the obtained plots, we additionally apply a linear interpolation and thereby generate a smooth landscape surface.
-Table 1: Summary of the FLA metrics used in our landscape analysis framework.
-Metrics
-Symbol Domain What a Higher Value Implies
+While it may seem intuitive that HP loss landscapes would differ depending on the target ML model, in practice the fact is often that common HPO methods could perform robustly under different ML systems.
+This inspires us to investigate whether the general family of HP loss landscapes may share certain inherent structural similarities.
 […]
-$\mathcal{L}$-ast
-$[-1,1]$
-HP Configurations with similar $\mathcal{L}$ values are more likely to be neighbors to each other.
+Our value proposition in this paper hinges on the hypothesis that with a deeper understanding of the underlying HP loss landscapes of ML systems, we can advance the interpretability and trustworthiness of the HPO process, and enhance the prospects of it as a tool to benefit broader scientific domains.
+2 HP Loss Landscape Construction and Analysis Method
+This section introduces our FLA framework developed for exploratory analysis of HP loss landscapes.
+We begin with a formal introduction of the HP loss landscape and how it can be constructed.
+We then provide a brief overview of our fine-grained landscape visualization method, selected classical FLA features, as well as our quantitative measures for landscape similarity assessment.
+Due to the stringent page limit, we only provide a sketch of these methods and more details are in Appendix B.
 […]
-$\rho_{a}$
-$[-1,1]$
-The landscape is smoother
-Neutrality Distance Correlation
-NDC
-$[-1,1]$
-The landscape is more likely to be flatter near the optimum.
+Landscape construction as directed graphs.
+Given the inherent neighborhood structure in HP loss landscapes, representing it as a graph is then a natural choice.
+In this model, each vertex is a configuration $\bm{\lambda}\in\bm{\Lambda}$, with the associated loss $\mathcal{L}(\bm{\lambda})$ as node attribute.
+Each $\bm{\lambda}^{\prime}\in\mathcal{N}(\bm{\lambda})$, is connected to $\bm{\lambda}$ via a directed edge, where the direction of each edge depends on the relative values of $\mathcal{L}(\bm{\lambda})$ and $\mathcal{L}(\bm{\lambda}^{\prime})$, always pointing to the fitter one.
+Such a graph-based model for the HP loss landscape inherently captures its neighborhood structure and loss dynamics.
+Once constructed, it then allows the implementation of many FLA methods via highly efficient graph traversal algorithms.
+Landscape topography plays a crucial role in governing the search behaviors of HPO algorithms.
+To quantitatively characterize its features, we pick up $6$ dedicated FLA metrics, each focusing on a specific aspect, and together they offer a comprehensive view of the essential landscape properties:
+  • Autocorrelation [28]: $\rho_{a}\in[-1,1]$, it measures the ruggedness of the landscape (i.e., degree of fluctuations in fitness values), where a higher value implies a smoother landscape surface.
+  • Loss assortativity, $\mathcal{L}_{\text{ast}}\in[-1,1]$: adapted from the general assortativity measure of complex networks [58], this shows the tendency that configurations with similar loss values are neighbors to each other, where a large value implies more local clustering.
+  • Mean neutrality, $\bar{\nu}\in[0,1]$: it is the Spearman correlation between the configurations neutralities and their distances to the global optimum; it quantifies the proportion of search space where fitness values remain constant (up to a tolerance $\epsilon=0.1\%$ in ML context) in local neighborhoods;
+  • Neutrality distance correlation, $\text{NDC}\in[-1,1]$: it assesses the tendency that the terrain becomes flatter when approaching the global optimum (i.e., diminishing returns).
+  • ##\##Local optima, $n_{\text{lo}}$, and mean basin size, $\bar{s}_{\mathcal{B}}$: local optima with a large basin of attraction can pose a greater challenge to optimizers, and so does a larger number of local optima.
+Landscape visualization is crucial for enabling an intuitive understanding of a landscape’s topography, but is notoriously difficult because it is often in a high-dimensional space.
+Some prior works address this by simply plotting only one or two HP(s) each time [41, 42], which fail to capture the global topography.
+Others works directly applied dimensionality reduction techniques to project the landscape into a $2$D space [38, 39, 40], but the results are not able to preserve the local neighborhood structure.
+We tackle this by leveraging a graph embedding method, HOPE [59], to first learn a vector representation for each configuration in the landscape, which can preserve high-order proximities between nodes.
+We then compress the obtained embeddings into $2$D using the UMAP [60] algorithm.
+To further refine interpretability, we additionally apply linear interpolation to generate a continuous landscape surface.
+We provide more details on design choices and configurations in Appendix B.1.
 […]
-$\bar{\nu}$
-$[0,1]$
-There are many ‘plateaus’ in the landscape.
-No. Local Optima
-$n_{\text{lo}}$
-$\mathbb{N}^{+}$
-There are many ‘valleys’ or ‘peaks’ in the landscape.
-Mean Basin Size
-$\bar{s}_{\mathcal{B}}$
-$\mathbb{R}^{+}$
-The local optima are hard to be escaped from.
- ${}^{1}$ewman (2010); ${}^{2}$einberger (1990), ${}^{3}$eidys & Stadler (2001)
-Quantifying Landscape Characteristics.
-To quantitatively assess the structural characteristics of HP loss landscapes, we employ a series of dedicated FLA metrics summarized in Table 1 as surrogate features.
-There are many other metrics for characterizing landscape properties (see Zou et al.
-(2022) for a detailed review), but our selected ones are particularly useful for this study as they cover the most essential landscape properties (i.e., modality, neutrality and smoothness) that are related to algorithm behaviors.
-More importantly, they are intuitive enough even for non-experts in FLA.
-Landscape Similarity in Terms of Performance Ranking.
-The comparison of rankings of HP configurations’ performance is the essence of a large corpora of HPO methods (Hutter et al., 2019).
-We thereby ground our landscape similarity measure of HP loss landscapes on the consistency of their performance ranks, denoted as $R(\mathcal{L}(\bm{\lambda}))$, to allow more informative results in the HPO context.
-Specifically, we use $3$ statistical metrics with complementary perspectives: 1)1)1 ) Spearman’s $\rho_{s}$, it measures the association of the performance ranks of configurations in two landscapes (Spearman, 1961), 2)2)2 ) Kaggle’s Shake-up metric (Trotman, 2019), it assesses the average movement of configuration rankings across two landscapes.
-3)3)3 ) The $\gamma$-set similarity (Watanabe et al., 2023a), it quantifies the ratio of overlaps between top-$10\%$ regions of two landscapes divided by their unions.
-In addition to these, to investigate the consistency of HP importance and interaction under different scenarios, We apply the widely used functional ANOVA method (Hutter et al., 2014a) to assess the variance contribution of every HP $\lambda\in\bm{\lambda}$ as well as their interactions up to the $3^{\text{rd}}$ order.
 3 Experimental Setup
-Table 2: Summarization of meta-information of our empirical study.
-  Models Datasets Fidelities Summarization
- Model
-Total
-HPs
-Total
-Configs.
-Cat.
-Class.
-Cat.
-Reg.
-Num.
-Class.
-Num.
-Reg.
-Image
-Class.
-Training
-Data
-Training
-Epochs
-Total
-Configs.
-# Land-
--scapes
-XGB $5$ $14,960$ $15$ $7$ $17$ $18$$\{0.1,0.25,\text{all}\}$ $2.56$M $342$
-RF $6$ $11,250$ $15$ $7$ $17$ $18$$\{0.1,0.25,\text{all}\}$ $1.92$M $342$
-LGBM $5$ $13,440$ $15$ $7$ $17$ $18$$\{0.1,0.25,\text{all}\}$ $2.30$M $342$
-DT $5$ $24,200$ $15$ $7$ $17$ $18$$\{0.1,0.25,\text{all}\}$ $4.14$M $342$
-CNN $8$ $6,480$ $6$$\{0.1,0.25,\text{all}\}$$\{10,25,50\}$ $0.35$M $108$
+Table 1: Summary of the models and landscape settings included in our empirical study.
+Group Model #HPs #Tasks #Fidelities #Configs. #Landscapes
+Classical DT $5$ $57$ $\alpha\in\{0.1,0.25,1\}$ $24,200$ $342$
+SVM $4$- $\alpha\in\{0.1,0.25,1\}$ $2,340$-
+Ensemble XGB $5$ $57$ $\alpha\in\{0.1,0.25,1\}$ $14,960$ $342$
+RF $6$ $57$ $\alpha\in\{0.1,0.25,1\}$ $11,250$ $342$
+LGBM $5$ $57$ $\alpha\in\{0.1,0.25,1\}$ $13,440$ $342$
+DNNs CNN $8$ $6$ $\alpha\in\{0.1,0.25,1\}$, $E\in\{10,25,50\}$ $6,480$ $108$
+NB101 NAS $1$ $E\in\{4,12,36,108\}$ $432,624$ $8$
+NB201 NAS $3$ $E\in\{10,50,200\}$ $15,625$ $18$
+NB360 NAS $2$ $E\in\{10,50,200\}$ $15,625$ $12$
+Investigating HPO scenarios posed in Section 1 demands the construction of HP loss landscapes on a wide spectrum of models, tasks and fidelities, and a vast amount of HP configurations.
+To this end, we conduct large-scale experiments to collect such data, as summarized in Table 1, which consumes over $7$ thousand of GPU and CPU hours (Appendix A) and ends up in $1,500+$ distinct landscapes.
 […]
-  Total (Before accounting $5$-fold cross-validation): $\bm{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}11.15\textbf{M}}}$ $\bm{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}1,500}}$
-Table 2 summarizes the meta-information of our empirical study, while detailed HP search space of each model and the principles we follow in designing them, are left in Appendix F.1.
-We first consider decision tree (DT(Safavian & Landgrebe, 1991) and three types of its ensembles: random forest (RF(Breiman, 2001), XGBoost (Chen & Guestrin, 2016) and LightGBM (Ke et al., 2017).
-We analyze the HP space of these models using the tabular benchmark proposed in Grinsztajn et al.
-(2022), which comprises $25$ regression and $32$ classification tasks (see Appendix F.2).
-These datasets span a broad range of complexities in terms of number of instances and features and are thus idea choice for comprehensive inspection of landscape characteristics.
-In addition to these, we also study convolutional neural networks (CNNs(Krizhevsky et al., 2012) on six classic image classification tasks (Appendix F.2) using a joint architecture and hyperparameter search (JAHS) (Bansal et al., 2022) space.
-We additionally consider another JAHS scenario, for which we adopt the NASBench-HPO (Klein & Hutter, 2019) data included in HPOBench (Eggensperger et al., 2021).
-This includes $62,208$ configurations of a feed-forward neural network (FCNet) evaluated on $4$ UCI datasets
-For each dataset, unless predefined, we randomly split the data into training ($80\%$) and test ($20\%$) set.
+HP search spaces.
+Due to the presence of continuous HPs, a full enumeration of many HP spaces is not possible.
+Instead, following [31], we adopt a grid sampling approach to construct HP loss landscapes.
+Note that while grid search is less efficient than methods like random search in terms of optimization [5], it can yield more interpretable landscapes (e.g., the neighboring configurations only differ in one HP) that are favorable by FLA.
+Also, though landscape extrapolation via surrogate models can offer infinite sampling resolution [17, 35], they introduce additional biases to the analysis [73].
+An important consideration in designing the grid spaces for each model is to ensure that they are representative of the settings commonly used in real-world practice, while keeping a good balance between grid resolution and computational cost.
+In doing so, we strictly followed our devised protocols in Appendix G.1 to create each space, with typically $2$ to $8$ HPs, and a few to tens of thousands of configurations (Table 1; Appendix G.1).
+These spaces are not meant to cover the entire feasible region, but rather to focus on more prominent and informative regions as in HPO.
+Datasets and configuration evaluation.
+For the first two groups of models, except SVM, we ground our analysis on $57$ tasks (Appendix G.2) from the tabular benchmark proposed in [67], which span a broad range of complexities (i.e., number of entries and features).
+For SVM, due to its poor scalability, we are only able to evaluate it on a few tasks.
+As for CNN, we consider $6$ image classification tasks as detailed in Appendix G.2.
+For each task, unless predefined, we randomly split the data into training ($80\%$) and test ($20\%$) sets.
 For all HP configurations $\bm{\lambda}\in\bm{\Lambda}$ of each model, we exhaustively evaluate $\mathcal{L}(\bm{\lambda})_{\text{train}}$ and $\mathcal{L}(\bm{\lambda})_{\text{test}}$ using $5$-fold cross-validation.
-Here, we use root mean squared error (RMSE) and $R^{2}$ score to serve as the loss function $\mathcal{L}$ for regression tasks, and for classification, we use accuracy and ROC-AUC.
-We control the fidelity of the training by varying the number of training instances to $\{10\%,25\%,100\%\}$ of the whole training data.
-For CNN, we additionally set the budget for the number of epochs to $\{10,25,50\}$ and thus obtain a total of $3\times 3$ different levels of fidelity.
-For FCNet, we vary fidelity by using meta-data at the $\{10,50,100\}$-th epoch.
-At the end, we obtain a total of $1,500$ landscapes with more than $11$M distinct HP configurations.
-To further demonstrate the transferability and potential impact of our proposed landscape analysis frameowork, we also employ it to analyze NASBench-101 (Ying et al., 2019), a well-known neural architecture search (NAS) benchmark.
+We use root mean squared error (RMSE) and $R^{2}$ score to serve as the loss function $\mathcal{L}$ for regression tasks, while for classification, we use accuracy and ROC-AUC.
+We control the fidelity of the training by either varying the ratio of training data ($\alpha$) or the number of training epochs ($E$, Table 1), thus obtaining cheaper proxies of the loss, $\mathcal{L}_{\text{LF}}$.
+Distances and neighborhood.
+After obtaining the set of HP configurations $\bm{\Lambda}$ and their corresponding losses $\mathcal{L}$ on each scenario, we then follow [31] to assess the distance between categorical HPs with the Hamming distance, where each differing value is assigned a $d=1$; for numerical HPs, we use the Manhattan distance, to reflect the number of discrete steps between them on the search grid.
+The aggregate distance between two configurations, $\bm{\lambda}_{i}$ and $\bm{\lambda}_{j}$, is the sum of these component distances, and we say they are neighbors to each other if $d(\bm{\lambda}_{j},\bm{\lambda}_{i})=1$.
+Such a neighborhood structure imitates the manual “baby-sitting” tuning process [13] and preserves the local proximity of configurations.
 4 Results and Analysis
-In this section, we seek to investigate HP loss landscapes under the four scenarios posed in Section 1.
-We will start from providing an universal view of the general characteristics of HP loss landscapes of different ML models (Section 4.1).
-We then explore and compare landscapes under: $i)$ training and test setups (Section 4.2), $ii)$ different fidelities (Section 4.3), $iii)$ different datasets (Section 4.4).
-Figure 2: 2D visualization of HP loss landscapes of $6$ ML models under different scenarios: (a) $\mathcal{L}_{\text{test}}$ landscape on baseline datasets (44059 for tree-based models, CIFAR-10 for CNN, protein structure for FCNet), (b) $\mathcal{L}_{\text{train}}$ landscape on baseline datasets, (c) Low-fidelity $\mathcal{L}_{\text{test}}$ landscape on baseline datasets, (d) $\mathcal{L}_{\text{test}}$ landscape on different datasets (44143 for tree-based models, Fashion MINIST for CNN, slice localization for FCNet).
-Colors indicate $R(\mathcal{L})$ (lower rank values are better).
-Figure 3: Distribution of FLA metrics introduced in Table 1 for each model across all datasets for landscapes of 1)1)1 ) $\mathcal{L}_{\text{test}}$, 2)2)2 ) $\mathcal{L}_{\text{train}}$ and 3)3)3 ) $\mathcal{L}_{\text{test}_{\text{LF}}}$.
-Figure 4: Distribution of Spearman, Shake-up and $\gamma$-set metrics between (a) $\mathcal{L}_{\text{test}}$ and $\mathcal{L}_{\text{train}}$, (b) $\mathcal{L}_{\text{test}}$ and $\mathcal{L}_{\text{test}_{\text{LF}}}$, (c) $\mathcal{L}_{\text{test}}$ across datasets.
-Medians are labeled beside each plot.
-4.1 Overall Characteristics of HP Loss Landscape of ML Models
-From landscape visualizations depicted in Figure 2 (a), we have a general impression that HP loss landscapes of ML models are highly structured and share certain patterns: they are relatively smooth; configurations are clustered in terms of performance; there is a highly distinguishable plateau consisting of prominent configurations, where the terrain becomes flatter.
-This impression is consistent with the FLA metrics reported in Figure 3, from which we see that landscapes for all models are:
-Fairly smooth and clustered.
-The high $\mathcal{L}$-ast and $\rho_{a}$ values for $\mathcal{L}_{\text{test}}$ landscapes shown in Figure 3 (a) and (b) respectively imply that configurations with similar $\mathcal{L}_{\text{test}}(\bm{\lambda})$ tend to be locally connected, where a small change in $\bm{\lambda}$ is not likely to cause dramatic variation of $\mathcal{L}_{\text{test}}(\bm{\lambda})$.
-This observation is similar to the findings in reinforcement learning (RL) (Eimer et al., 2023), where the transitions between different parts of the HP landscapes of RL are also found to be quite smooth.
-This property makes the HP landscapes favorable to Bayesian optimization and search space pruning techniques, as it would be easier to separate the promising regions from the poorly performing ones.
-On the other hand, if the landscape is rugged instead, in which $\mathcal{L}_{\text{test}}(\bm{\lambda})$ of very different levels often mixed together, it would become more difficult for such techniques to identify a clear promising region.
+We now investigate the obtained HP loss landscapes under the $4$ scenarios posed in Section 1.
+We begin with exploring the general characteristics of these landscapes in Section 4.1.
+We then compare landscapes under: $i)$ training and test setups (Section 4.2), $ii)$ different fidelities (Section 4.3), and $iii)$ different datasets (Section 4.4).
+Note that as we only constructed a few landscapes for SVM, we left its results in Appendix C, where our findings here also generalize to SVM.
+Figure 2: $2$D visualization of HP loss landscapes using our proposed method in Section 2 for (A) CNN and (B) XGBoost, w.r.t.
+$\mathcal{L}_{\mathrm{test}}$ and $\mathcal{L}_{\mathrm{train}}$ as well as different datasets and fidelities (by varying the fraction of training data $\alpha$ or the number of training epochs $E$, marked in red color).
+(C) shows the visualization of the NASBench-101 landscape with different training epochs on CIFAR-10, and (D) visualizes the NASBench-201 (❶–❸) and NASBench-360 (❹–❺) loss landscapes of different datasets.
+Colors indicate ranks of performance (lower rank values are better) to normalize the losses.
+Figure 3: Distribution of FLA metrics across all datasets for landscapes of 1)1)1 ) $\mathcal{L}_{\mathrm{test}}$, 2)2)2 ) $\mathcal{L}_{\mathrm{train}}$ and 3)3)3 ) $\mathcal{L}_{\mathrm{test}}$ obtained with $10\%$ training data.
+4.1 Overall Characteristics of Studied HP Loss Landscapes
+We first visualize the topography of the $\mathcal{L}_{\mathrm{test}}$ landscapes for each model using the method in Section 2.
+Fig.
+2 (A-1) and (B-1) show the results for CNN and XGBoost.
+These provide a general impression of these landscapes, and we clearly observed certain structural similarities: the HP loss landscapes of the studied models are relatively smooth; configurations are locally clustered in terms of their losses; there is a highly distinguishable plateau consisting of prominent configurations, where the terrain becomes flatter.
+As depicted in Appendix D.1, the landscapes of the other models also exhibit similar patterns.
+We will then take a closer look at these patterns with quantitative FLA metrics in Fig.
+3.
+Remark 1. We further verified the reliability of these visualizations in Appendix D.2 with tunable $NK$-landscapes [74] and the black-box optimization benchmark functions from COCO platform [75].
+Remark 2. The patterns observed for the systems in Fig. 2 are representative of the other systems, as justified quantitatively by the landscape metrics in Fig. 3 and landscape similarities in Fig. 4.
+Fairly smooth and locally clustered.
+From Fig.
+3, we found high loss assortativity and autocorrelation values for the majority of our studied $\mathcal{L}_{\mathrm{test}}$ landscapes.
+These imply that configurations with similar $\mathcal{L}_{\mathrm{test}}(\bm{\lambda})$ tend to be neighbors to each other, where a small change in $\bm{\lambda}$ is not likely to cause drastic variation in $\mathcal{L}_{\mathrm{test}}(\bm{\lambda})$ (i.e., a fairly smooth landscape).
+Such property is highly beneficial, as it makes an elegant separation of the prominent regions from the poorly performing ones possible.
+Such separation can be known as priors, and be utilized to guide the design of search spaces that focus on the promising region [76], or to warm-start HPO solvers directly within that region [53].
+Alternatively, this knowledge can also be learned on-the-fly as the optimization proceeds, and thus guides the solver towards fitter areas [4], or assisted in progressively pruning the search space to narrow down the search [77, 78].
+On the contrary, if the landscape is highly rugged instead, in which prominent configurations are often interspersed with poorer ones, it would become more difficult to distinguish a clear advantageous region, making it more difficult to navigate the landscape or design a compact search space.
 Nearly unimodal.
-As shown in Figure 3 (e), we find that a considerable fraction of the $\mathcal{L}_{\text{test}}$ landscapes are unimodal, whereas for other landscapes, there could be a handful to dozens (DT) of local optima.
-This is somewhat contradictory to Pushak & Hoos (2022) at the first thought, in which the authors found that almost all landscapes they studied are nearly unimodal.
-However, when taking a closer look at the local optima in our landscapes, we find that they usually feature a small basin of attraction (Figure 3 (f)).
-This makes them relatively ‘shallow’, and thus would not pose significant obstacles to optimization.
-However, beyond the results in Figure 3, we find that FCNet landscapes on the $4$ UCI datasets possess $24$ to $347$ local optima, with $\bar{s}_{\mathcal{B}}$ up to $2,372$ (Appendix D), implying a strong attraction for optimizers.
-Pushak & Hoos have also reported similar observations on these four landscapes, and they speculated the reason could be that these scenarios fall into the over-parameterized regime.
-While we agree with this reasoning, we seek to conduct further analysis on the local optima using the local optima network (LON) (Ochoa et al.
-(2008), Appendix B.3).
-We find that despite the pressence of many other local optima, the global optima still plays a pivotal role in the connectivity of the LON (Appendix D).
-Therefore, many local optima can eventually escape the global optimum via cetain strategies (e.g., a perturbation), though this may take additional efforts.
+As shown in Fig.
+3 (E), we found a considerable fraction of the $\mathcal{L}_{\mathrm{test}}$ landscapes are unimodal, while for some landscapes (DT, especially), there could be a handful to dozens of local optima.
+This is somewhat contradictory to [31] at first glance, in which it is found that almost all landscapes studied are nearly unimodal.
+However, when taking a closer look at the local optima in our landscapes, we found that they are usually featured in a small basin of attraction (see Fig.
+3 (F)).
+This makes them relatively “shallow”, and would not pose significant obstacles to optimization.
+Thus, we contend that the unimodal hypothesis is still valid for most of our studied scenarios.
 Highly neutral; planar around the optimum.
-As depicted in Figure 3 (d), we can clearly see that HP loss landscapes are often featured in high neutrality.
-This indicates that a large portion of $1$-bit moves in the landscape will result in subtle change in $\mathcal{L}_{\text{test}}(\bm{\lambda})$ (i.e., $\leq 1\text{\textperthousand}$).
-We postulate a major reason for this is the low effective dimensionality (Bergstra & Bengio, 2012) of HPO problems: usually only a small subset of all available HPs have obvious influence on performance.
-Despite landscape neutrality can largely vary with the choice on which HPs to analyze and their respective values, considering the fact that we have ready removed totally unimportant HPs from teh search space, moves with subtle performance shifts can actually be more prevalent than one may expect.
-Such phenomenon is more pronounced for the well-performing regions, as illustrated by the high NDC values in Figure 3 (c).
-It suggests that as we move closer to the global optimum, it is more likely to encounter neutral moves, and the landscape becomes flatter.
-This is in line with Probst & Boulesteix (2017); Pimenta et al.
-(2020) and practical experience: the gain of tuning HPs usually progressively decreases as approaching the best reachable performance.
-Such property also poses challenges to optimizers, as there is little gradient information that can be utilized for navigation towards fitter configurations (Muñoz et al., 2015).
-Overall, despite certain exceptions, we see that the family of HP loss landscapes tend to share various high-level properties, whereas the exact topologies would vary with models.
-This explains why in practice, one can often expect an optimizer to work relatively robustly across a wide range of scenarios.
-In addition, most properties here also seem to generalize the NAS problems (Appendix C), except that we find the NAS landscapes tend to have lower neutrality and more local optima.
-4.2 Training and Test HPO Landscapes
-Figure 2 (a) and (b) provide a visual comparison of $\mathcal{L}_{\text{train}}$ and $\mathcal{L}_{\text{test}}$ landscapes for different models, respectively.
-We could infer from the plots that the structural characteristics of the $\mathcal{L}_{\text{train}}$ landscapes highly our previously discussed properties.
-On the other hand, for the performance rankings, we notice that $\mathcal{L}_{\text{train}}$ generally correlate with $\mathcal{L}_{\text{test}}$ for RF, DT, LGBM and CNN, whereas for XGBoost, there is significant shifts in performance between the two cases.
-We further quantitatively verify such observations using the FLA metrics and the landscape similarity metrics introduced in Section 2.
-Structural characteristics.
-From Figure 3, we could clearly see that the landscape characteristics for $\mathcal{L}_{\text{train}}$ and $\mathcal{L}_{\text{test}}$ are highly consistent for most studied scenarios.
-More specifically, it is surprising to see that $\mathcal{L}_{\text{train}}$ landscapes tend to yield relatively higher $\mathcal{L}$-ast and $\rho_{a}$, suggesting a smoother and more structured terrain.
+As in Fig.
+3 (D), we clearly observed that HP loss landscapes are often featured in high neutrality.
+This indicates that a large portion of purturbations in the landscapes will result in subtle changes in $\mathcal{L}_{\mathrm{test}}$ (i.e., $\leq 1\text{\textperthousand}$).
+We postulate that a major reason for this is the low effective dimensionality [5] of HPO problems, i.e., usually only a small subset of all tunable HPs have a significant impact on performance.
+As our designed search spaces have excluded HPs that have no obvious impact on performance, the high neutralities of the resulting landscapes imply that neutral plateaus can actually be more prevalent than one may expect.
+This phenomenon is more pronounced for the well-performing regions, as illustrated by the high NDC values in Fig.
+3 (C).
+It suggests that as we move closer to the global optimum, it is more likely to encounter neutral moves, and the landscape thus becomes flatter.
+This observation coincides with experimental observations in HPO, i.e., the gain of tuning HPs often progressively diminishes as approaching higher performance.
+Such property can pose challenges to the solvers, as there is little gradient information that can be utilized for navigating towards fitter configurations, and they can ‘get lost’ in neutral plateaus.
+This then underscores the necessity of careful search space design prior to optimization.
+Remark 3.
+The exact measured neutrality of a landscape can depend on multiple factors, including the choice of HPs, the domain, and resolution of the search space, as well as the tolerance $\epsilon$ we used to define neutral moves.
+Though we have made careful considerations in ensuring the generalizability of our results (see Appendix G), there are no golden standards or consensus for us to explicitly follow.
+Our goal here is to evoke attention to the potential redundancy in search spaces used in HPO practice, rather than obtaining an absolute measure of neutrality in these landscapes.
+The NAS landscapes, as evidenced in Appendix E, also yield high assortativity and autocorrelation values, and the terrain tends to be flatter near the optimum as well.
+These properties are not surprising as both NAS and HPO are sibling subjects under the umbrella of AutoML.
+Albeit this, we found that NAS landscapes exhibit considerably more local optima, constituting a multimodal topography (see, e.g., Fig.
+2 C-5).
+In this regard, the NAS problems seem to be more complex.
+Therefore, we can expect that directly applying HPO solvers to NAS may not lead to satisfactory performance.
+Instead, tailored search strategies are required to tackle such more rugged landscape surfaces.
+Figure 4: Distribution of Spearman, Shake-up and $\gamma$-set metrics between (A) $\mathcal{L}_{\mathrm{test}}$ and $\mathcal{L}_{\mathrm{train}}$, (B) $\mathcal{L}_{\mathrm{test}}$ and $\mathcal{L}_{\mathrm{test}}$ with $10\%$ training data, (C) $\mathcal{L}_{\mathrm{test}}$ across datasets.
+4.2 HP Loss Landscapes on Training and Test Data
+Fig.
+2 (A-2), (B-2) visualize the $\mathcal{L}_{\mathrm{train}}$ landscapes of CNN and XGBoost (see also Appendix D.1).
+The structural characteristics of these landscapes show high concordance with our previously discussed properties for $\mathcal{L}_{\mathrm{test}}$.
+On the other hand, for performance rankings, we observed that the $\mathcal{L}_{\mathrm{train}}$ generally correlate with $\mathcal{L}_{\mathrm{test}}$ for CNN, RF, DT, and LGBM, whereas for XGBoost, there is significant discrepancy between the landscapes of $\mathcal{L}_{\mathrm{train}}$ and $\mathcal{L}_{\mathrm{test}}$.
+We further quantitatively discuss such observations in the following paragraphs.
+Landscape characteristics.
+From Fig.
+3, we can clearly see that the structural characteristics for $\mathcal{L}_{\mathrm{train}}$ and $\mathcal{L}_{\mathrm{test}}$ landscapes are highly consistent for most studied scenarios.
+We found that $\mathcal{L}_{\mathrm{train}}$ landscapes tend to yield relatively higher performance assortativity and autocorrelation values and thus a smoother and more structured terrain.
 Meanwhile, the NDC values are lower in the training scenarios.
-These observations imply that $\mathcal{L}_{\text{train}}$ landscapes are even more benign than $\mathcal{L}_{\text{test}}$ landscapes.
-In addition, we find that $\bar{\nu}$, $n_{\text{lo}}$ and $\bar{s}_{\mathcal{B}}$ values rarely change between $\mathcal{L}_{\text{train}}$ and $\mathcal{L}_{\text{test}}$ landscapes.
-Notably, local optima found in $\mathcal{L}_{\text{train}}$ and $\mathcal{L}_{\text{test}}$ landscapes are almost (if not all) identical.
-These indicate that the relative performance in local neighborhoods tend to remain similar in the two cases, despite the variations in their numerical values and the global performance rankings.
-Landscape similarity in terms of performance rankings.
-We quantified the similarity between all pairs of $\mathcal{L}_{\text{train}}$ and $\mathcal{L}_{\text{test}}$ landscapes for all $5$ models using the three metrics introduced in Section 2 as shown in Figure 4 (a).
-Overall, we observe that $R(\mathcal{L}_{\text{train}})$ and $R(\mathcal{L}_{\text{test}})$ are globally correlated for all models except XGBoost, as indicated by the significant $\rho_{s}$ values (median $>0.7$) and low Shake-up metrics (median $<0.15$).
-However, when zooming into the top-$10\%$ regions, we find that the majority of our studied scenarios reveal low $\gamma$-set similarities.
+These observations imply that the $\mathcal{L}_{\mathrm{train}}$ landscapes are even more benign than $\mathcal{L}_{\mathrm{test}}$ ones.
+In addition, we found that neutrality, #local optima, and basin sizes rarely change between $\mathcal{L}_{\mathrm{train}}$ and $\mathcal{L}_{\mathrm{test}}$ landscapes.
+Notably, #local optima found in $\mathcal{L}_{\mathrm{train}}$ and $\mathcal{L}_{\mathrm{test}}$ landscapes are almost (if not all) identical.
+This suggests that the relative performance in local neighborhoods tends to be retained in the two cases, despite variations in their numerical values and global rankings.
+Landscape similarity.
+We quantified the similarity between all pairs of $\mathcal{L}_{\mathrm{train}}$ and $\mathcal{L}_{\mathrm{test}}$ landscapes using the $3$ metrics introduced in Section 2 as shown in Fig.
+4 (A).
+Overall, we observed that $R(\mathcal{L}_{\mathrm{train}})$ and $R(\mathcal{L}_{\mathrm{test}})$ are globally correlated for all models except XGBoost, as indicated by the significant $\rho_{s}$ values (median $>0.7$) and low Shake-up metrics (median $<0.15$).
+However, when zooming into the top-$10\%$ regions, we saw that the majority of our studied scenarios reveal low $\gamma$-set similarities.
 It indicates that the generalization gap is larger in prominent regions where configurations are highly adapted to the training set.
-This phenomenon is more severe for XGBoost, where the median $\gamma$-set similarity is only $0.07$, and there is also a poor $\rho_{s}$ value (median = $0.34$) and high Shake-up score (median = $0.25$).
-In order to gain more insight into such generalization gaps for XGBoost, we create scatter plots of $\mathcal{L}_{\text{test}}$ versus $\mathcal{L}_{\text{train}}$ on dataset #44059 as shown in Figure 5 (a).
-We decompose the pattern into two modes: During the first mode, $\mathcal{L}_{\text{test}}$ highly correlates with $\mathcal{L}_{\text{train}}$ as it decreases, and the models in this stage underfit the data.
-In the next mode, as points struggle to further move on the $x$-axis ($\mathcal{L}_{\text{train}}$), they stagnate or even significantly increase on the $y$-axis ($\mathcal{L}_{\text{test}}$), indicating strong evidence of overfitting.
-In particular, we can see a plateauing trend near the $x$-axis, where some models overly excel on the training data, but performing poorly on the test set.
-To further investigate which kinds of configurations are likely to lead to overfitting, we color the points with respect to their HP values as shown in Figure 5 (b-e).
-We are excited to see that the generated plots demonstrate clear patterns between the value of each HP and the resulted performance.
-In particular, we find that learning rate, max depth and subsample have significant impact on $\Delta\mathcal{L}$.
+This is more severe for XGBoost, where the median $\gamma$-set similarity is only $0.07$, and there is also a low $\rho_{s}$ value (median = $0.34$) and high Shake-up score (median = $0.25$).
+A closer look at overfitting.
+In order to gain more insight into such generalization gaps for XGBoost, we create scatter plots of $\mathcal{L}_{\mathrm{test}}$ versus $\mathcal{L}_{\mathrm{train}}$ on the dataset #44059 in Fig.
+5 (A), where we can decompose the pattern into two modes.
+In the first one, $\mathcal{L}_{\mathrm{test}}$ highly correlates with $\mathcal{L}_{\mathrm{train}}$ as it decreases, and the models in this stage underfit the data.
+In the second mode, on the contrary, as points struggle to further move on the $x$-axis ($\mathcal{L}_{\mathrm{train}}$), they stagnate or even significantly increase on the $y$-axis ($\mathcal{L}_{\mathrm{test}}$), indicating strong evidence of overfitting.
+In particular, we can see a plateauing trend near the $x$-axis, where some models overly excel on the training data, but perform poorly on the test set.
+To further investigate which kinds of configurations are likely to lead to overfitting, we color the points with respect to their HP values as shown in Fig.
+5 (B-E).
+It is interesting to see that the generated plots demonstrate clear patterns between the value of each HP and the resulting performance.
+In particular, we find that learning rate, max depth, and subsample have a significant impact on $\Delta\mathcal{L}$.
 However, the generalizability of a learner is not monopolized by a single one of them; instead, it depends on their cumulative interactions.
-For example, the largest $\Delta\mathcal{L}$s are observed for learners that features a large learning rate, deep base trees, combined with low subsample rate, but any of these HP settings alone does not necessarily lead to the worst case performance.
-In addition to this, we notice that such generalization gap is also related to dataset characteristics and weakly correlated across models, and we dissuss more about this matter in Appendix E.
-Figure 5: (a) Scatter plot of $\mathcal{L}_{\text{train}}$ versus $\mathcal{L}_{\text{test}}$ for all $14,960$ configurations of XGBoost on the dataset #44059.
-$\bm{\lambda}^{*}_{\text{test}}$ is marked by red star , (b-e) The same plot with colormap added according to HPs values.
+For example, the largest $\Delta\mathcal{L}$s are observed for learners that feature a large learning rate, and deep base trees, combined with a low subsample rate, but any of these HP settings alone does not necessarily lead to the worst case performance.
+In addition to this, we noticed that such a generalization gap is also correlated some dataset properties such as number of samples or features, and we discuss more about this in Appendix F.
+Figure 5: (A) Scatter plot of $\mathcal{L}_{\mathrm{train}}$ versus $\mathcal{L}_{\mathrm{test}}$ of XGBoost on the dataset #44059.
+$\bm{\lambda}^{*}_{\mathrm{test}}$ is marked by , (B-E) The same plot with colors to indicate HPs values.
 Warmer color indicate higher values.
-4.3 HPO Landscapes with Different Fidelities
-Figure 2 (c) shows the low-fidelity test loss landscapes (denoted as $\mathcal{L}_{\text{test}_{\text{LF}}}$) for each model (using $10$ epochs for FCNet, and $10\%$ training data for others).
-From the plots, we could see that $\mathcal{L}_{\text{test}_{\text{LF}}}$ landscapes are highly consistent with $\mathcal{L}_{\text{test}}$ in terms of both structural characteristics and performance rankings.
-More specifically, as reflected in Figure 3, all measured FLA metrics of $\mathcal{L}_{\text{test}_{\text{LF}}}$ landscapes showed little difference compared to $\mathcal{L}_{\text{test}}$ landscapes across all studied scenarios.
-For performance rankings, Figure 4 (b) depicts the distribution of the $3$ similarity indicators between $\mathcal{L}_{\text{test}_{\text{LF}}}$ and $\mathcal{L}_{\text{test}}$ across all datasets for each model.
-We could observe a high Spearman correlation (median $>0.85$) between $\mathcal{L}_{\text{test}}$ and $\mathcal{L}_{\text{test}_{\text{LF}}}$ for all models, and the $\gamma$-set similarities between the top-$10\%$ configurations are also prominent, with medians larger than $60\%$.
-These imply that $R(\mathcal{L}_{\text{test}})$ and $R(\mathcal{L}_{\text{test}_{\text{LF}}})$ are highly consistent for the majority of our studied scenarios and there is large overlap between the promising regions of the two landscapes.
-In addition, the Shake-up scores yield low values ($\text{median}<0.1$), suggesting that on average the difference between $R(\mathcal{L}_{\text{test}})$ and $R(\mathcal{L}_{\text{test}_{\text{LF}}})$ is less than $10\%$.
-Additional results on FCNet, NAS-Bench-101 in Appendix D and Appendix C respectively are also consistent with our findings here.
-4.4 HPO Landscapes Across Datasets
+4.3 HP Loss Landscapes with Different Fidelities
+Fig.
+2 (A-3) and (B-3) show the low-fidelity loss landscapes (denoted as $\mathcal{L}_{\mathrm{test}}^{\mathrm{LF}}$) for CNN and XGBoost with $\alpha=0.1$ training data.
+From the plots, we observed that $\mathcal{L}_{\mathrm{test}}^{\mathrm{LF}}$ landscapes are highly consistent with $\mathcal{L}_{\mathrm{test}}$ in terms of both structural characteristics and performance rankings.
+More specifically, as reflected in Fig.
+3, all measured FLA metrics of $\mathcal{L}_{\mathrm{test}}^{\mathrm{LF}}$ landscapes exhibited little difference compared to $\mathcal{L}_{\mathrm{test}}$ landscapes across most studied scenarios.
+For performance rankings, Fig.
+4 (B) depicts the distribution of the $3$ similarity indicators between $\mathcal{L}_{\mathrm{test}}^{\mathrm{LF}}$ and $\mathcal{L}_{\mathrm{test}}$ across all datasets for each model.
+We observed a high Spearman correlation (median $>0.85$) between $\mathcal{L}_{\mathrm{test}}$ and $\mathcal{L}_{\mathrm{test}}^{\mathrm{LF}}$ for most cases, and the $\gamma$-set similarities between the top-$10\%$ configurations are also prominent, with medians larger than $60\%$.
+These imply that $R(\mathcal{L}_{\mathrm{test}})$ and $R(\mathcal{L}_{\mathrm{test}}^{\mathrm{LF}})$ are highly consistent for the majority of our studied scenarios and there is a large overlap between the promising regions of the two landscapes.
+In addition, the Shake-up scores yield low values ($\text{median}<0.1$), suggesting that on average the difference between $R(\mathcal{L}_{\mathrm{test}})$ and $R(\mathcal{L}_{\mathrm{test}}^{\mathrm{LF}})$ is less than $10\%$.
+Note that albeit our general observations, the similarity distributions in Fig.
+4 (B) typically have long tails, and there are cases for which the $\mathcal{L}_{\mathrm{test}}^{\mathrm{LF}}$ landscapes largely deviate from the ground truth.
+Figure 6: $\rho_{s}$ between $\mathcal{L}_{\mathrm{test}}^{\textrm{LF}}$ and $\mathcal{L}_{\mathrm{test}}$ in NASBench-201 as a function of training epoch.
+The NAS landscapes also suggest similar findings.
+In Fig.
+2 (C), we plot the NASBench-101 $\mathcal{L}_{\mathrm{test}}^{\mathrm{LF}}$ landscapes under four different training epochs on CIFAR-10.
+While it is not surprising that with higher fidelities (i.e., $E$), the $\mathcal{L}_{\mathrm{test}}^{\mathrm{LF}}$ landscapes can better approximate the ground truth $\mathcal{L}_{\mathrm{test}}$ landscape, we noticed that with only $4$ training epochs, the $\mathcal{L}_{\mathrm{test}}^{\mathrm{LF}}$ landscapes can already capture certain prominent regions.
+We further explored this with NASBench-201, which provides evaluation information after every training epoch.
+Fig.
+6 plots the Spearman correlation between $\mathcal{L}_{\mathrm{test}}^{\mathrm{LF}}$ and $\mathcal{L}_{\mathrm{test}}$ as a function of training epochs $E$.
+From this we found that with $10$ out of $200$ training epochs, we can already obtain a reasonable ranking.
+4.4 HP Loss Landscapes Across Datasets
+Figure 7: Importance of top-$5$ HPs (or interactions) as indicated by functional ANOVA method of each model.
+Performance rankings.
+Fig.
+2 (A-4) and (B-4) show the $\mathcal{L}_{\mathrm{test}}$ landscapes for CNN and XGBoost on different datasets.
+We can see that the high-level topography of the $\mathcal{L}_{\mathrm{test}}$ landscape are preserved when transferring to a new task.
+It is interesting to note that the top regions in the original landscape generally retain their positions, despite the changes in their exact contours.
+The FLA metrics reported in Fig.
+3 also support such observation.
+In addition, from the similarity metrics reported in Fig.
+4, we observed that the measured performance reveal clear Spearman correlations (median $>0.65$) across datasets.
+In practice, stakeholders are more concerned about the top performing regions in the landscape.
+For this, the overlap between the top-$10\%$ regions, as indicated by the $\gamma$-set similarity, also achieves medians around $40\%$.
+We also found that the landscape similarities for tabular tasks are correlated with simple dataset meta-features (see Fig.
+15 in Appendix F), and there have been works that study such features in more depth (e.g., [79]).
+HP importance and interaction.
+In addition to performance rankings, Fig.
+7 illustrates the contribution of each HP and their interactions to model performance assessed by the functional ANOVA method.
+The results indicate that some (or combination terms of) HPs are typically important for many datasets for a given model.
+For example, the depth-conv contributes a large portion of variance to model performance for CNN, and so does its interactions with width-conv.
+Such knowledge can also be utilized to expedite the HPO process.
+Our additional results for NAS-Bench-201 across CIFAR-10/100 as well as ImageNet-16-120 datasets also revealed similar highly-transferable conclusions (see Fig.
+2 D1-D3).
+However, again, we note that our findings are based on general trends, and there are outliers in Fig.
+4 that reveal little landscape similarity between tasks.
+Also, the results on NASBench-360 (see Fig.
+2 D4-D5), which evaluates the search space of NASBench-201 on two tasks very different from classic image classification datasets, exhibited considerable deviations from the landscapes in Fig.
+2 D1-D3.
+Considering that these two tasks are rooted in biomechanics and computational physics respectively, the resulting difference might be reasonable.
+However, this evokes the need for including a more diverse set of tasks to further validate our findings, and we leave this for future works.
+5 Discussions and Conclusions
+Many studies have significantly advanced the methodological developments of HPO.
+Yet, the understanding of the underlying HP loss landscapes of ML models as well as their correlations, which are the foundations of HPO, has been largely overlooked.
+By analyzing $1,500+$ HP loss landscapes of $6$ representative ML models induced on different scenarios with fine-grained visualizations and dedicated FLA metrics, we found a considerable resemblance between test loss landscapes and their training counterparts, albeit the choice of certain HP combinations and datasets can lead to significant discrepancies.
+We also showed that the topographies of HP loss landscapes with lower fidelities, in general, are highly consistent with the full-fidelity landscapes, and the same applies to landscapes of the same model induced on different datasets.
+We further revealed a unified portrait of the HP loss landscapes across our studied models in terms of their structural characteristics.
+For example, they tend to be fairly smooth, nearly unimodal, and featured with a high degree of neutrality.
+Our results in part reaffirm isolated patterns observed in prior works [31, 32, 33, 34, 35] regarding landscape modality and neutrality, and extend them to broader scenarios.
+At the same time, we provide the first landscape-level evidence to support the effectiveness of transfer learning and multi-fidelity methods, which previously have been mainly based on rough intuition and empirical performance.
+Moreover, the high-level structural similarities shared by our studied HP loss landscapes induced on different models assist in explaining why many HPO solvers often work robustly across different systems.
+Also, given the smooth topography of these landscapes, we also highlight the importance of elaborated search space design, including both the selection of HPs and the choice of value domains.
 […]
-Figure 6: Importance of top-$5$ HPs as determined by functional ANOVA method for each model.
-In addition to performance rankings, Figure 6 illustrates the contribution of each HP and their interactions to model performance assessed by the functional ANOVA method.
-The results indicate that some (combination of) HPs are typically important for many datasets for a given model.
-For example, learning rate consistently contributes a large portion of variance to model performance for LightGBM, and its interactions with the number of leaves and estimators are also important.
-These observations are similar with van Rijn & Hutter (2018), which find also conclude that certain HPs of a ML model are important for a wide spectrum of datasets by analyzing meta-data from OpenML platform.
-As discussed in Section 4.1, HP loss landscapes often involve a large number of none-improvement moves, especially near the optimum.
-We also see that there is clear division between the promising regions and the poorly-performaning ones.
-Therefore, leveraging prior knowledge from previously tasks should be able to greatly expedite the searching process by means like warm-starting HPO from good configurations or carefully selecting candidate HPs and crafting the search space.
-More importantly, based on our results, we note that this should not be only limited to similar tasks defined under certain rules, since they may not always available.
-On the other hand, seemingly different tasks could still provide informative information as we see above.
-Our additional results for FCNet on $4$ datasets, and NAS-Bench-201 across CIFAR-10/100 as well as ImageNet datasets (Appendix C), also revealed similar highly-transferable conclusions.
-5 Disuccsions and Conclusions
-By conducting large-scale exploratory analysis on $1,500$ HP landscapes of $6$ ML models with over $11$M model configurations under different scenarios, this paper reveals an unified portrait of their topographies in terms of smoothness, neutrality and modality.
-We also show that these properties are highly transferable across datasets and fidelities, and thus provide fundamental evidence to support the effectiveness of transfer and multi-fidelity methods, which in privious practices, is mainly based on intuition.
-However, while our findings are observed for the majority studied scenarios, we do observe some exceptions.
-For example, most landscapes inspected reveal a nearly unimodal structure, but some of them can have dozens to a few hundreds of local optima with non-negligible basin sizes (e.g., FCNet).
-Also, there are cases when landscapes under lower fidelities or on a different task reveal very different patterns, as shown by the long tails of the similarity distributions in Figure 4.
-Further explorations by interrogating the relationship between dataset characteristics may provide a even more comprehensive understanding of the HPO landscape.
-Our developed FLA framework in this work, has shown great potential for enabling both qualitative and quantitative understanding for a wider range of AutoML problems.
-While it currently relies on large-scale, pre-evaluated data points for landscape construction, we think it can be a promising direction to integrate it with existing AutoML frameworks, and thus allow for on-the-fly analysis of problem landscapes, thereby allowing it to be accessible by a broader range of stakeholders.
-Appendix A Background and Related Work
-Fitness Landscape of BBOPs.
-A central topic in optimization concerns how the intricate interplay of decision variables determine the targeted objective values.
-For example, geneticists and biologists are interested in the relationship between genotypes and the phenotypes and functions of organisms that may affect their evolutionary success (De Visser & Krug, 2014).
-In chemical physics, scientists seek to explore how the energy of a system is dependent on its various configurational or conformational states (Brooks III et al., 2001).
-The fitness landscape can be envisioned as a surface over the high-dimensional space as formed by decision variables, where the objective values are represented as the elevation of the landscape.
-The structure of this surface describes the spectrum of possible objectives values across the variable space and thus strongly influences the optimization.
-In addition, searching trajectories of problem solvers can be thought as strategic walks on the corresponding problem landscape.
-Successful applications of this metaphor to analyze black-box systems have advanced the understanding of many fields (He & Liu, 2016; Puchta et al., 2016; Shires & Pickard, 2021; Brooks III et al., 2001).
-Most related to our work is the fitness landscape analysis (FLA) for black-box optimization problems (BBOPs) (Malan, 2021).
-Many metrics that quantify the structural characteristics of fitness landscapes have been developed over the years to describe the topography of BBOP landscapes, e.g., autocorrelation, neutrality, modality, epistasis variance, evolvability, etc.
-Fitness Landscape on other ML Systems.
-In addition to the works that focus on analyzing HP and AutoML landscapes introduced in Section 1, there are also works which applied FLA to study other ML systems, including the neural architecture search (NAS) spaces (Rodrigues et al., 2020; Traoré et al., 2021) and the HP landscapes of reinforcement learning (RL) (Eimer et al., 2023; Mohan et al., 2023).
-We note that these works are somewhat orthogonal to ours, as we focus on the HP landscapes of ML algorithms.
-In addition, there is also another line of work on investigating the loss landscapes of neural networks (Rakitianskaia et al., 2016; Rodrigues et al., 2022; van Aardt et al., 2017).
-These works, though seemingly similar, are focused on the neural network losses which are the objectives of weight optimizers during model training, and are thus significantly different from ours.
-HPO Landscape Visualization.
-The ability to visualize HP loss landscapes is crucial for enabling intuitive understandings of their complex topologies, which is however, far from trivial in practice due to their high-dimensional nature.
-Current study methods usually bypass this issue by exmaining only two HPs each time (see, e.g., Akiba et al.
-(2019); Ørebæk & Geitle (2021)).
-However, such techqniues can not provide an unified illustration of the total interactive effects of all interested HPs.
-In the evolutionary computation, there are have been successful attempts on applying dimensionality reduction techniques to visualize the fitness landscape of BBOPs (Michalak, 2019; Huang & Li, 2023; Sass et al., 2022; Walter et al., 2022).
-Similar work has also been done in the domain of chemical physics (Shires & Pickard, 2021), where UMAP and $t$-SNE are used to visualize the so-called energy landscape.
-Overfitting in Machine Learning.
-To empirically investigate the presence of overfitting in real-world ML applications, Roelofs et al.
-(2019) analyzed massive submission meta-data on the Kaggle platform, and their results suggest little evidence of overfitting due to testset reuse.
-However, since the Kaggle data is constituted by a diverse set of ML models with very different HP configurations that are not comparable with each other, it prohibits further analysis on which HP configurations are more likely to lead to overfitting.
-Hyperparameter Importance and Interaction.
-Understanding which HPs influence model performance to what extend can provide valuable insights into the tuning strategy (Probst et al., 2019).
-To analyze importance of hyperparameters, one could either use models that are inherently interpretable, e.g., decision trees (Quinlan, 1986), or apply model-agnostic methods such as functional ANOVA (Hutter et al., 2014a; Watanabe et al., 2023b).
-Built on this technique, van Rijn & Hutter (2017) compared the performance of a variety of algorithms on large set of OpenML datasets and stated that the same hyperparameters were typically important across datasets.
-Appendix B Details of the Landscape Analysis Framework
-Figure 7: High-level conceptual workflow of our proposed HP loss landscape analysis framework.
-Figure 7 illustrates the conceptual workflow of our proposed HP loss landscapes analysis framework in Section 2.
-From a high level, our landscape visualizations (Appendix B.1) seek to provide a general sketch of the landscape topography.
-The observed patterns can then be quantitatively verified via a set of dedicated FLA metrics characterizing complementary aspects of landscape properties (Appendix B.2), while landscapes visualizations, at the same time, can assist the intuitive interpretation of the numerical figures.
-In addition to these, we also apply $3$ dedicated similarity measures to quantify the consistency of configuration performance across landscapes (Appendix B.4).
+Despite these caveats, our exploratory study may provide a useful perspective to rethink HPO algorithms by considering the topography of the underlying HP loss landscapes upon which they traverse.
+There could also be nuanced scrutiny of the scenarios discussed in this study, e.g., the relationship between dataset meta-features and the landscape characteristics.
+Also, our developed FLA framework could serve as a useful lens for future studies on a broader range of search landscapes in the ML community (e.g., NAS, AutoRL, and AutoML), and beyond (e.g., protein fitness landscapes [?] toward AI for science).
+We hope our analysis opens a new avenue towards understanding the fitness landscapes of black-box systems and explainable AI in optimization.
+Appendix A Summary of Computational Resrouces Used
+[…]
+[…]
+Appendix B Details of the FLA Framework
 B.1 Landscape Visualization Method
@@
 node embedding method to serve this purpose, as it is able to capture asymmetric high-order proximity in directed networks.
-Specifically, in a directed network [^2] Most manipulations and analyses of HP loss landscapes as complex networks are grounded on NetworkX and Pandas package., if there is a directed link from vertex $v_{i}$ to vertex $v_{j}$ and from vertex $v_{j}$ to vertex $v_{k}$, it is more likely to have a link from $v_{i}$ to $v_{k}$, but not from $v_{k}$ to $v_{i}$.
+Specifically, for a directed network [^2] Most manipulations and analyses of HP loss landscapes as complex networks are grounded on NetworkX and Pandas package., if there is a directed edge from vertex $v_{i}$ to vertex $v_{j}$ and from vertex $v_{j}$ to vertex $v_{k}$, it is more likely to have a edge from $v_{i}$ to $v_{k}$, but not from $v_{k}$ to $v_{i}$.
 In order to preserve such asymmetric transitivity, HOPE learns two vertex embedding vectors $U^{s},U^{t}\in\mathbf{R}^{|V|\times d}$, which is called source and target embedding vectors, respectively.
-After constructing the high-order proximity matrix $S$ from $4s$ proximity measures, i.e., Katz Index, Rooted PageRank, Common Neighbors and AdamicAdar.
+After constructing the high-order proximity matrix $S$ from $4$ proximity measures, i.e., Katz Index, Rooted PageRank, Common Neighbors and AdamicAdar.
 HOPE finally learns vertex embeddings by solving the a matrix factorization problem:
@@
 While HOPE (as well as any other node embedding methods) could generate vectorized embeddings for configurations, these embeddings typically are still high-dimensional and not compatible for $2$D visualization.
-To cope with this, we further apply UMAP [^3] We use the implementation in UMAP package.
+In light of this, we further apply UMAP [^3] We use the implementation in UMAP package.
 to project the HOPE embeddings into a $2$D space, which is based on three assumptions, and here we provide brief discussion on whether they hold for our case:
-  • Data are uniformly distributed on Riemannian manifold.
+[…]
+    Data is uniformly distributed on Riemannian manifold.
 As the distribution of HOPE embeddings essentially depends on the connectivity pattern of the graph, for HP loss landscapes, we can expect this assumption to hold.
 This is because the number of neighbors for each node in the graph would largely stay the same based on our neighborhood definition.
-Therefore, there will be no significance variation of density across the graph, and the HOPE embeddings in turn should approximately have a uniform distribution.
+Therefore, there will be no significance variation in density across the graph, and the HOPE embeddings in turn should approximately have a uniform distribution.
+[…]
     The Riemannian metric is locally constant.
 The distances between HOPE embeddings correlate directly with the local connectivity patterns within the network.
-Given the stability in neighborhood structures as we discussed above, it is reasonable to presume that the local distance metrics would remain approximately constant in small regions.
+Given the stability in neighborhood structures as we discussed above, it is reasonable to presume that local distance metrics would remain approximately constant in small regions.
+[…]
     The manifold is locally connected.
@@
 We thereby expect this assumption will also hold.
-Remark on Algorithm Choices.
+Remark on algorithmic choices.
 In principle, other node embedding and dimensionality reduction methods can be applied to serve our purpose.
 Our specific choice on HOPE is mainly because of its scalability to large-scale networks and its ability to preserve both local and global structure of the landscape.
-As for dimensionality reduction, Draganov et al.
-(2023) has made detailed theoretical comparisons between UMAP and $t$-SNE, and shows that only the normalization significantly impacts their outputs.
-This thens implies that a majority of the algorithmic differences can be toggled without affecting the embeddings.
+As for dimensionality reduction, [83] has made detailed theoretical comparisons between UMAP and $t$-SNE, and shows that only the normalization significantly impacts their outputs.
+This then implies that a majority of the algorithmic differences can be toggled without affecting the embeddings.
 We choose UMAP here for its better scalability.
@@
 Since HP loss landscapes can vary a lot in terms of dimensionality and total number of configurations, there is no ‘one-size-fits-all’ setup for HPs of UMAP.
-However, we still provide some general guidelines for HP tuning, which mainly focus on two most important HPs, namely n_neighbors and min_dist.
+Yet, here we provide some empirical guidelines for HP tuning, which mainly focus on two most important HPs, namely n_neighbors and min_dist.
 Specifically, n_neighbors controls the balance between local and global structure in the embedding.
 In general, we found that it is better to set n_neighbors to a value that is larger than the average number of neighbors for each node.
-For min_dist, which specifies the minimum distance between points in the low-dimensional space, we generally recommend values larger than $0.5$.
+For min_dist, which specifies the minimum distance between points in the low-dimensional space, we recommend values larger than $0.5$.
 The reasoning here is that we want the points to be more spread out in the low-dimensional space, and thus prevent the points to be densely distributed in local regions.
@@
 B.2 Landscape Analysis Metrics
+0:  A starting configuration $\mathbf{c}$; A neighborhood function $\mathcal{N}_{d}$; A fitness function $f$ 1:  while $\mathbf{c}$ is not a local optimum do 2:    $\mathbf{c}^{\prime}_{\text{best}}=\arg\min_{\mathbf{c}^{\prime}\in\mathcal{N}_{1}(\mathbf{c})}f(\mathbf{c}^{\prime})$ 3:    if $f(\mathbf{c}^{\prime}_{\text{best}})<f(\mathbf{c})$ then 4:       $\mathbf{c}\leftarrow\mathbf{c}^{\prime}_{\text{best}}$ 5:    end if 6:  end while 7:  return $\mathbf{c}$
 Algorithm 1 Best-Improvement Local Search
-0:  A starting configuration $\mathbf{c}$; A neighborhood function $\mathcal{N}_{d}$; A fitness function $f$
-1:  while $\mathbf{c}$ is not a local optimum do
-2:     $\mathbf{c}^{\prime}_{\text{best}}=\arg\min_{\mathbf{c}^{\prime}\in\mathcal{N}_{1}(\mathbf{c})}f(\mathbf{c}^{\prime})$
-3:     if $f(\mathbf{c}^{\prime}_{\text{best}})<f(\mathbf{c})$ then
-4:        $\mathbf{c}\leftarrow\mathbf{c}^{\prime}_{\text{best}}$
-5:     end if
-6:  end while
-7:  return $\mathbf{c}$
+0:  The set of all configurations $\mathcal{C}$ 1:  $\mathcal{V}\leftarrow\emptyset$ 2:  $\mathcal{B}\leftarrow\emptyset$ 3:  for all $\mathbf{c}\in\mathcal{C}$ do 4:    $\mathbf{c}^{\ell}\leftarrow\textsc{LocalSearch}(\mathbf{c})$ 5:    $\mathcal{B}[\mathbf{c}^{\ell}]\leftarrow\mathcal{B}[\mathbf{c}^{\ell}]\cup\{\mathbf{c}\}$ 6:  end for 7:  return $\mathcal{V},\mathcal{B}$
 Algorithm 2 Identifying Local Optima and Their Basins
-0:  The set of all configurations $\mathcal{C}$
-1:  $\mathcal{V}\leftarrow\emptyset$
-2:  $\mathcal{B}\leftarrow\emptyset$
-3:  for all $\mathbf{c}\in\mathcal{C}$ do
-4:     $\mathbf{c}^{\ell}\leftarrow\textsc{LocalSearch}(\mathbf{c})$
-5:     $\mathcal{B}[\mathbf{c}^{\ell}]\leftarrow\mathcal{B}[\mathbf{c}^{\ell}]\cup\{\mathbf{c}\}$
-6:  end for
-7:  return $\mathcal{V},\mathcal{B}$
 Assortativity coefficient.
@@
 For HP loss landscape, the performance assortativity measures the extent to which HP configurations with similar performance are more likely to be neighbors to each other.
-Formallly, given a HP loss landscape as a directed graph, and the model loss $\mathcal{L}$ takes values $[L_{1},L_{2},\dots]$, the $\mathcal{L}$-assortativity evaluates the Pearson correlation of the measured loss between pairs of linked configurations and is measured as (Newman, 2003):
--$\mathcal{L}\text{-ast}=\frac{\sum_{i}e_{ii}-\sum_{i}a_{i}b_{i}}{1-\sum_{i}a_{i}b_{i}}$ (2)
-where $e_{ij}$ is called mixing matrix entry, which represents the fraction of total edges in the network (i.e., landscape) which connects configurations having performance $\mathcal{L}(\bm{\lambda})=L_{i}$ to configurations having attribute $\mathcal{L}(\bm{\lambda})=L_{j}$.
+Formallly, given a HP loss landscape as a directed graph, and the model loss $\mathcal{L}$ takes values $[L_{1},L_{2},\dots]$, the $\mathcal{L}$-assortativity evaluates the Pearson correlation of the measured loss between pairs of linked configurations and is measured as [58]:
+$\mathcal{L}_{\text{ast}}=\frac{\sum_{i}e_{ii}-\sum_{i}a_{i}b_{i}}{1-\sum_{i}a_{i}b_{i}}$ (2)
+where $e_{ij}$ is called mixing matrix entry, which represents the fraction of total edges in the network (i.e., landscape) that connects configurations having performance $\mathcal{L}(\bm{\lambda})=L_{i}$ to configurations having attribute $\mathcal{L}(\bm{\lambda})=L_{j}$.
 In directed networks like our case, this can be asymmetric, i.e., $e_{ij}\neq e_{ji}$.
 In addition, $a_{i}=\sum_{j}e_{ij}$ is the portion of edges $(\bm{\lambda}_{u},\bm{\lambda}_{v})$ such that $\mathcal{L}(\bm{\lambda}_{u})=L_{i}$ and $b_{i}=\sum_{j}e_{ji}$ is the portion of edges $(\bm{\lambda}_{v},\bm{\lambda}_{u})$ such that $\mathcal{L}(\bm{\lambda}_{v})=L_{i}$.
-A high $\mathcal{L}$-ast would imply that configurations with similar performance have strong tendancy to be connected and form local clusters.
+A high $\mathcal{L}_{\text{ast}}$ would imply that configurations with similar performance have strong tendancy to be connected and form local clusters.
 Landscape neutrality.
 It is often in genetics that a mutation in a single position of a DNA sequence will only lead to negligible change in the expression.
-Such phenomenon is known as landscape neutrality at a macro level, and for each sequence, this can be quantitatively measured by the mutational robustness (Payne & Wagner, 2019), which is the probability for such non-effective mutation to happen among all its possible mutants.
+Such phenomenon is known as landscape neutrality at a macro level, and for each sequence, this can be quantitatively measured by the mutational robustness [84], which is the probability for such non-effective mutation to happen among all its possible mutants.
 Similar ideas are also applicable to HPO, where altering certain HPs may only result in subtle performance shifts.
@@
 We quantitatively assess this using the neutrality distance correlation (NDC), which measures the Pearson correlation coefficient between the neutrality of a configuration $\bm{\lambda}$ and its distance to the global optimum, $\mathrm{d}(\bm{\lambda},\bm{\lambda}^{*})$ (equation (4)).
-We need to note that result derived from this can be potentially affected by the choice of $\epsilon$ for neutrality.
-To cope with this, we also develop an alternative method for assessing NDC, which is directly based on raw loss differences between configurations.
 Specifically, for each adaptive walk in the landscape using best-improvement local search (Algorithm 1) that can eventually approach $\bm{\lambda}^{*}$, we measure the Pearson correlation coefficient between $\Delta\mathcal{L}$ for each pair of consecutive configurations $(\bm{\lambda}_{i},\bm{\lambda}_{i-1})$ ($i\geq 2$), and $\mathrm{d}(\bm{\lambda}_{i},\bm{\lambda}^{*})$.
 We calculate NDC as the average across all such walks.
-However, in this paper, to keep the consistency with the neutrality metric, we report the results based on the first method.
 $\text{NDC}=\rho_{p}[\nu(\bm{\lambda}),\mathrm{d}(\bm{\lambda},\bm{\lambda}^{*})]$ (4)
@@
 For a unimodal landscape, there is only a global optimum configuration $\bm{\lambda}^{*}$.
-In constrast, multimodal landscapes have various local optima with sub-optimal performance, which can pose challenges to the optimization.
+In constrast, multimodal landscapes have various local optima with sub-optimal performance, and is known to be ‘rugged’.
 Size of basin of attraction.
@@
 For each landscape, we conduct $100$ random walks of length $100$ and average $\rho_{a}$ across all measurements to mitigate the effects of randomness.
-B.3 Local Optima Network
-Beyond the number of local optima and their basin sizes, a further aspect to investigate is the connectivity pattern between them.
-For example, an important question that we may concern is whether we can escape from a given local optimum to the global optimum?, if yes, then what is the chance of this?
-Local optima networks (LON) (Ochoa et al., 2008; Vérel et al., 2011), which are rooted in the study of energy landscapes in chemical physics (Stillinger, 1995), address these questions by constructing a subspace of the original fitness landscape where the nodes indicate local optima, and edges represent possible transitions between them.
+B.3 Local Optima Network (LON)
+Beyond the number of local optima and their basin sizes, a further landscape property to interrogate is the distribution of local optima and the connectivity pattern between them.
+For example, an important question that we may concern is whether we can escape from a given local optimum to the global optimum?, if yes, then, what is the chance of this?
+Local optima networks (LON) [29, 85], which are rooted in the study of energy landscapes in chemical physics [86], address these questions by constructing a subspace of the original fitness landscape where the nodes indicate local optima, and edges represent possible transitions between them.
 In particular, an improving edge can be traced from local optimum configuration $\bm{\lambda}^{\ell}_{i}$ to $\bm{\lambda}^{\ell}_{j}$ if configurations in $\mathcal{B}(\bm{\lambda}^{\ell}_{i})$ can escape to $\bm{\lambda}^{\ell}_{j}$ by applying a $2$-bit perturbation followed by local search.
 The edge weights $w_{i,j}$ indicate the total probability for such transitions to happen (see Algorithm 3) between two local optima.
-By conducting network mining on LONs (e.g., Huang & Li (2023)), we can get further insights into the distribution and connection between local optima as well as how they are potentially linked with the global optimum.
-However, while we believe LON could be an effective tool for analyzing complex landscapes and we incorporate this as part of our landscape analysis tool, we do not present such analyses in most scenarios in this paper since their HP loss landscapes only possess a few local optima (if not none).
+By conducting network mining on LONs (e.g., [21]), we can get further insights into the distribution and connection between local optima as well as how they are potentially linked with the global optimum.
+0:  The set of local optima $\mathcal{V}$; The basin of attraction of each local optimum $\mathcal{B}_{\mathbf{c}^{\ell}}$; The set of all configurations $\mathcal{C}$; A neighborhood function $\mathcal{N}_{d}(\mathbf{c})$ 1:  $\mathcal{E}\leftarrow\emptyset$ 2:  $\mathcal{W}\leftarrow\emptyset$ 3:  for all $\mathbf{c}^{\ell}\in\mathcal{V}$ do 4:    for all $\mathbf{c}^{\ell\prime}\in\mathcal{N}_{2}(\mathbf{c}^{\ell})$ do 5:       $\mathbf{c}^{\ell}_{\text{new}}\leftarrow\textsc{LocalSearch}(\mathbf{c}^{\ell\prime})$ 6:       if $f(\mathbf{c}^{\ell}_{\text{new}})<f(\mathbf{c}^{\ell})$ then 7:          if edge $(\mathbf{c}^{\ell},\mathbf{c}^{\ell}_{\text{new}})$ not in $\mathcal{E}$ then 8:             $\mathcal{E}\leftarrow\mathcal{E}\cup\{(\mathbf{c}^{\ell},\mathbf{c}^{\ell}_{\text{new}})\}$ 9:             $\mathcal{W}[(\mathbf{c}^{\ell},\mathbf{c}^{\ell}_{\text{new}})]\leftarrow 1$ 10:          else 11:             $\mathcal{W}[(\mathbf{c}^{\ell},\mathbf{c}^{\ell}_{\text{new}})]\leftarrow\mathcal{W}[(\mathbf{c}^{\ell},\mathbf{c}^{\ell}_{\text{new}})]+1$ 12:          end if 13:       end if 14:    end for 15:  end for 16:  return $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{W})$
 Algorithm 3 Constructing Local Optima Network
-0:  The set of local optima $\mathcal{V}$; The basin of attraction of each local optimum $\mathcal{B}_{\mathbf{c}^{\ell}}$; The set of all configurations $\mathcal{C}$; A neighborhood function $\mathcal{N}_{d}(\mathbf{c})$
-1:  $\mathcal{E}\leftarrow\emptyset$
-2:  $\mathcal{W}\leftarrow\emptyset$
-3:  for all $\mathbf{c}^{\ell}\in\mathcal{V}$ do
-4:     for all $\mathbf{c}^{\ell\prime}\in\mathcal{N}_{2}(\mathbf{c}^{\ell})$ do
-5:        $\mathbf{c}^{\ell}_{\text{new}}\leftarrow\textsc{LocalSearch}(\mathbf{c}^{\ell\prime})$
-6:        if $f(\mathbf{c}^{\ell}_{\text{new}})<f(\mathbf{c}^{\ell})$ then
-7:           if edge $(\mathbf{c}^{\ell},\mathbf{c}^{\ell}_{\text{new}})$ not in $\mathcal{E}$ then
-8:              $\mathcal{E}\leftarrow\mathcal{E}\cup\{(\mathbf{c}^{\ell},\mathbf{c}^{\ell}_{\text{new}})\}$
-9:              $\mathcal{W}[(\mathbf{c}^{\ell},\mathbf{c}^{\ell}_{\text{new}})]\leftarrow 1$
-10:           else
-11:              $\mathcal{W}[(\mathbf{c}^{\ell},\mathbf{c}^{\ell}_{\text{new}})]\leftarrow\mathcal{W}[(\mathbf{c}^{\ell},\mathbf{c}^{\ell}_{\text{new}})]+1$
-12:           end if
-13:        end if
-14:     end for
-15:  end for
-16:  return $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{W})$
 B.4 Landscape Similarity Metrics
@@
 Shake-up metric.
-It originates from the Kaggle competition community and is designed to assess the rank changes between the public and private leaderboards of a competition (Trotman, 2019).
+It originates from the Kaggle competition community and is designed to assess the rank changes between the public and private leaderboards of a competition [63].
 To be specific, it quantifies the average movement of rankings from public board to the private board.
@@
 $\gamma$-set similarity.
-It is proposed in Watanabe et al.
-(2023a) to assess the similarity of two tasks using the ratio of their most proninent configurations.
+It is proposed in [64] to assess the similarity of two tasks using the ratio of their most proninent configurations.
 More specifically, for two HP loss landscapes, their $\gamma$-set similarity is defined as the ratio of the intersection of the top-$\gamma$ regions to the union of them.
 In this paper, we consider $\gamma=10\%$.
-Appendix C Analysis on NAS Loss Landscapes
-C.1 NAS Benchmarks
-NAS-Bench-101.
+Appendix C HP Loss Landscapes of SVM
+Figure 8: HP loss landscapes of SVM on four tabular datasets, with $\alpha=\{0.1,1\}$ training data, with respect to training and test accuracy.
+Due to the inherent scalability issue of SVM, it is infeasible to evaluate it on most of our considered tabular datasets, as they often contains a large sample size and a high-dimensional feature space.
+Instead, we only conducted experiments on a few datasets that are relatively small in size with respective to two HPs (gamma and c) using RBF kernel.
+Since the resulting HP loss landscapes are already in a 2D space, we thus directly visualize them as heatmaps.
+Fig.
+8 presents the results on four tabular datasets (#44122, #44123, #44089, #44125) with $\alpha=\{0.1,1\}$ training data.
+We can observe that the patterns here generally match our previous findings on other models, i.e., the landscapes are smooth, highly clustered, and nearly unimodal; a considerable portion of the landscape is neutral, and the performance difference diminishes around the optimum.
+These general characteristics also largely preserves across datasets, or with lower training fidelities.
+Nevertheless, similar to our observed patterns for XGBoost, here the training loss landscape exhibits very different topography compared to the test loss landscape, where overfitting is prevalent.
+Appendix D Additional Results on Landscape Visualization.
+D.1 Landscape Visualizations for DT, LGBM and RF
+We present additional landscape visualizations for decision tree (DT), LightGBM (LGBM), and random forest (RF) in Fig. 9.
+Figure 9: 2D visualization of HP loss landscapes of DT, LightGBM and RF under different scenarios: (A) $\mathcal{L}_{\text{test}}$ landscape on baseline datasets (#44059), (B) $\mathcal{L}_{\text{train}}$ landscape on baseline datasets, (C) Low-fidelity $\mathcal{L}_{\text{test}}$ landscape on baseline datasets obtained using $10\%$ training data, (D) $\mathcal{L}_{\text{test}}$ landscape on different datasets (#44143).
+Colors indicate ranks of performance (lower rank values are better).
+D.2 Landscape Visualizations for NK-Landscape and BBOB Functions
+Figure 10: Visualizations of Kauffman’s $NK$-landscape of $n=10$ with different levels of ruggedness using our proposed method in Section 2 (low: $k=0$, medium: $k=5$, and strong: $k=10$).
+Figure 11: Recovered landscapes with our proposed visualization method versus groundtruth landscape (reference) surfaces of $f_{1}$ and $f_{17}$ to $f_{24}$ in the black-box optimization benchmark (BBOB) benchmarking suite from COCO.
+In order to demonstrate the capability of our proposed landscape visualization method in capturing the structural characteristics of fitness landscapes, here we conduct additional experiments with artifical black-box optimization problems.
+NK-landscape.
+We first consider the well-known $NK$-landscape model [74], which was introduced by Stuart Kauffman in the last century for studying the evolutionary of biological systems.
+It has two tunable parameters: $n$ and $k$.
+Here, $n$ represents the number of elements in the system, while $k$ determines the level of interaction between these elements.
+As $k$ increases, the landscape becomes increasingly rugged, with a greater number of local optima.
+For illustration here, we generate $NK$-landscapes with $n=10$ and three levels of ruggedness by varying the parameter $k\in\{0,5,10\}$.
+As shown in Fig.
+10, our visualization method is able to capture the increasing ruggedness of the landscapes as $k$ grows, where the landscape becomes more rugged.
+BBOB benchmark suite.
+Another scenario we consider here is the black-box optimization benchmark (BBOB) functions suite from the COCO platform [75].
+This suite contains $24$ BBOB functions ($f_{1}$ to $f_{24}$) with a diverse range of complexities that are widely studied in literature.
+Here we test the capability of our visualization method in recovering the topography of these functions by comparing the visualized landscapes with the groundtruth surfaces.
+In particular, we present the results for $f_{17}$ to $f_{24}$ in Fig.
+11, which are known to exhibit strong multimodality (to compare, we also present $f_{1}$, a highly separable unimodal function).
+As can be seen from the results, our method is able to accurately recover shape of the function surface, even for the most complicated ones.
+Appendix E Analysis on NAS Loss Landscapes
+E.1 NAS Benchmarks
+NASBench-101.
 It represents the first effort towards benchmarking NAS research and thus fostering reproducibility in the community.
 It evaluates a CNN architecture with $3$ stacked blocks, where a down-sampling is added between two consecutive blocks; A $3\times 3$ convolution is used before the main blocks, and the outputs of the main blocks are fed to an average pooling and fully connected layer.
-NAS-Bench-101 considers a cell-based search space which enumerates all possible configurations for each block.
+NASBench-101 considers a cell-based search space which enumerates all possible configurations for each block.
 More specifically, the search space is formulated as a directed acylic graph (DAG) with $7$ nodes and a maximum of $9$ edges.
@@
 After removing all isomorphic cells, this search space results in $423$k unique configurations.
-NAS-Bench-101 evaluates each of them on the CIFAR-10 dataset and records meta-data at the $\{4,12,36,108\}^{\text{th}}$ epoch.
-NAS-Bench-201.
-It features a different skeleton compared to NAS-Bench-101, in which a residual block is applied to connect $3$ cells.
+NASBench-101 evaluates each of them on the CIFAR-10 dataset and records meta-data at the $\{4,12,36,108\}^{\text{th}}$ epoch.
+NASBench-201.
+It features a different skeleton compared to NASBench-101, in which a residual block is applied to connect $3$ cells.
 Each cell here is a DAG with $4$ nodes and $6$ edges.
@@
 The benchmark thus contains $5^{6}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}15,625}$ unique model architectures, with each evaluated on $3$ different datasets: $i)$: CIFAR-10, $ii)$: CIFAR-100, $iii)$: ImageNet-16-120 using $200$ epochs.
-C.2 NAS Landscape Construction
+[…]
+E.2 NAS Landscape Construction
 Despite the cell-based search spaces of NAS benchmarks are very different from the HP ones considered in this paper, our landscape construction rountine could be easily transfered to NAS by redefining the neighborhood structure.
 This then demands proper encoding of the NAS configurations and the definition of a suitable distance function.
-NAS-Bench-101 Neighborhood.
+NASBench-101 Neighborhood.
 In the original paper, a cell-encoding method based on adjacency matrices is introduced to encode configurations, which comprises two components.
@@
 The benchmark API provides built-in function to check for this.
-NAS-Bench-201 Neighborhood.
-The encoding of configurations in this benchmark is much more straight forward.
+NASBench-201/360 Neighborhood.
+The encoding of configurations in this search space is much more straight forward.
 Specifically, we encode each configuration using a $6$-bit vector, where each bit specifies the operation taken at each edge, which can take $5$ categorical values.
 We base our neighborhood definition here on the $1$-edit distance as well, in which a $1$-bit mutant of a configuration is the one with the operation in only one edge altered.
-C.3 Results on NAS Loss Landscapes
+E.3 Results on NAS Loss Landscapes
 While we have conducted all analyses on both benchmarks, it would be too tedious to lay all the information here.
-Instead, we use NAS-Bench-101 to discuss general NAS landscape characteristics and multi-fidelity, and present the comparisons between NAS landscapes across datasets using NAS-Bench-201.
-Before we present our results, we also note that while the majority of configurations ($359$k out of $423$k) in NAS-Bench-101 search space come with $7$ nodes (i.e., with $5$ intermediate operations), the rest of them have less nodes.
+Instead, we use NASBench-101 to discuss general NAS landscape characteristics and multi-fidelity, and present the comparisons between NAS landscapes across datasets using NASBench-201.
+Before we present our results, we also note that while the majority of configurations ($359$k out of $423$k) in NASBench-101 search space come with $7$ nodes (i.e., with $5$ intermediate operations), the rest of them have less nodes.
 Here we mainly focus on configuragtions with $7$ nodes, since accounting all the configurations would result in many independent components in the landscape.
@@
 NAS Landscape Visualization.
-We first visulize the NAS-Bench-101 landscape using our proposed landscape visualization method in Section 2, as shown in Figure 8.
+We first visulize the NASBench-101 landscape using our proposed landscape visualization method in Section 2, as shown in Fig.
+2 (C).
 Specifically, we plot the training accuracy landscape along with test accuracy landscapes trained at $4$ different number of epochs.
@@
 Considering that the search space here contains nearly $30$ times more confiurations than the HP space we used in the main text, each cluster may contain thousands of configurations, inside which the landscape could still be sufficiently smooth.
-Figure 8: Visualization of NAS-Bench-101 landscape using our proposed method.
-Here, we present the training accuracy landscapes as well as test accuracy landscapes at different recorded epochs.
-Color indicates rank of test accuracy, and higher values are better.
-NAS Landscape Metrics. Here we report several landscape metrics for the NAS-Bench-101 landscape with respect to test accuracy at the $108$-th epoch:
+NAS Landscape Metrics. Here we report several landscape metrics for the NASBench-101 landscape with respect to test accuracy at the $108$-th epoch:
   • Autocorrelation.
@@
   • Number of Local Optima.
-One of the most distinguishable property of NAS-Bench-101 is a large collection of local optima in the landscape.
+One of the most distinguishable property of NASBench-101 is a large collection of local optima in the landscape.
 In fact, we found $5,908$ local optima in total (out of the $359$k configurations), which could probably make the landscape far more difficult to optimization.
 We will discuss more about them and their basins in the LON part.
-Figure 9: Visualization of the local optima network of NAS-Bench-101 test landscapes at the $108$-th epoch, containing $5,908$ nodes.
+Figure 12: Visualization of the local optima network of NASBench-101 test landscapes at the $108$-th epoch, containing $5,908$ nodes.
 Radius of each node (i.e., local optimum) indicates size of the corresponding basin of attraction.
@@
 NAS Landscape Across Fidelities.
-From Figure 8, we could see that in general the test landscapes with lower fidelities resemble the one trained with $108$ epoch.
+From Fig.
+2 (C), we could see that in general the test landscapes with lower fidelities resemble the one trained with $108$ epoch.
 Quantitatively, the Spearman correlation between the test accuracy landscapes at $108$-th epoch and $32$-th epoch is $0.904$, with a Shake-up metric of $9.74\%$.
@@
 In general, the landscapes are still moderately correlated, but the detailed pattern can be largely distorted.
-Figure 10: Visualization of NAS-Bench-201 landscape using our proposed method across datasets.
+Figure 13: Distribution of local optima test accuracy versus LON metrics.
 Local Optima Network Analysis.
-Figure 11: Distribution of local optima test accuracy versus LON metrics.
-Figure 9 shows the local optima network of NAS-Bench-101 test landscapes at the $108$-th epoch.
+Fig.
+12 shows the local optima network of NASBench-101 test landscapes at the $108$-th epoch.
 It could be obviously seen that there are clear community (clustering) structure in the network, where local optima with large basin of attractions tend to locate at the center of each cluster, which usually feature a promising performance.
-To be specific, from the left plots in Figure 11, we could clearly observe that the size and radius [^5] Here, we define the radius of a basin of attraction as the expected number of local search steps needed to reach the corresponding local optimum.
+To be specific, from the left plots in Fig.
+13, we could clearly observe that the size and radius [^5] Here, we define the radius of a basin of attraction as the expected number of local search steps needed to reach the corresponding local optimum.
 Since in practice, we observe that this value is often highly correlated with the size of basin of attraction, and correlate with performance in a similar manner, we only report basin size in the main text.
 of the basin of attraction is positively correlated with the performance of local optima (Spearman correlation $>0.55$).
-More importantly, as suggested by the cumulative distribution of basin size shown in Figure 12, the highly-fit local optima have a dominant size of basin of attractions in the landscape.
-For example, the dashed line in Figure 12 indicates that the cumulative sum of basin sizes of those local optima with $acc>94.3\%$ take $50\%$ of the total basin size (which is the number of total configurations in the landscape).
+More importantly, as suggested by the cumulative distribution of basin size shown in Fig.
+14, the highly-fit local optima have a dominant size of basin of attractions in the landscape.
+For example, the dashed line in Fig.
+14 indicates that the cumulative sum of basin sizes of those local optima with $acc>94.3\%$ take $50\%$ of the total basin size (which is the number of total configurations in the landscape).
 Since being in the basin of a local optimum would imply that local search will eventually converge it, our result is to say that if we start local search from a random configuration in the landscape, there is $50\%$ chance that we would end up in a local optimum with $acc>94.3\%$.
@@
 Beyond local search, we then continue to consider if we do fall in a local optima whose fitness is not that satisfactory, then what is the chance that we can somehow escape from it?.
-The two plots at the right panel of Figure 11 show the distribution of escape rate and improve rate correlate with test accuracy.
+The two plots at the right panel of Fig.
+13 show the distribution of escape rate and improve rate correlate with test accuracy.
 Here, the escape rate is the chance that, a configuration in the basin of a local optimum $\mathbf{c}^{\ell}$, after applying a $2$-bit perturbation, would converge to a different local optimum $\mathbf{c}^{\ell}_{\text{new}}$.
 On the other hand, the improve rate further restricts that the new local optimum should have better performance compared to the current one.
-From Figure 11, we could clearly see observe the majority of the local optima feature a escape rate of larger than $50\%$, which suggests that most local optima are not that difficult to escape from.
+From Fig.
+13, we could clearly see observe the majority of the local optima feature a escape rate of larger than $50\%$, which suggests that most local optima are not that difficult to escape from.
 For improve rate, we observe a good correlation with test accuracy, where it is more easy to find an improving move for a poorly-performed local optimum.
 Unfortunately, for those that already have promising performance, there is only little chance to transit to a better basin using $2$-bit perturbation.
-C.4 Results on NAS-Bench-201
-We visualize the loss landscapes of NAS-Bench-201 on the $3$ datasets, namely CIFAR-10/100 and ImageNet, in Figure 10.
-In general, we see that results on these $3$ tasks reveal strong consistency to each other (Spearman correlation $>0.95$), which conforms with our findings on HP loss landscapes.
-Appendix D Details About FCNet HP Loss Landscapes
-Figure 13: Local optima network for the $100$ local optima in the FCNet landscape on parkinsons telemonitoring dataset.
-The left plot shows the full-view of the network, where node size represents the size of the corresponding basin of attraction.
-Node color indicates the performance of the model, which is labeled as rank values for each node, and the global optimum has a rank of $1$.
-The right plot shows the neighborhoods of the global optimum, while other nodes are omitted.
-Table 3: Landscape Features of FCNet on Different Datasets
- Dataset Autocorr. Assortativity Neutrality NDC #LO Mean Basin Size
-Parkinsons Tele. $0.4683$ $0.6232$ $0.4317$ $0.6368$ $100$ $589$
-Protein Struc. $0.4896$ $0.5789$ $0.4746$ $0.7411$ $128$ $385$
-Naval Prop. $0.5720$ $0.5298$ $0.4091$ $0.6433$ $347$ $159$
-Slice Local. $0.4887$ $0.6134$ $0.4427$ $0.8216$ $24$ $2372$
-D.1 Landscape Features
-Figure 14: Spearman correlation between LON features ($i)$: mean basin size, $ii)$: mean basin radius, $iii)$: maximum basin radius, $iv)$: in-coming strengths, $v)$: out-going strengths, $vi)$: closeness centrality, $vii)$: eigenvector centrality, $viii)$: pagerank centrality) and local optima performance for FCNet across datasets.
-From the visualization of FCNet landscapes shown in Figure 2, we can conclude that they generally follow the same characteristics as we discussed in Section 4.1.
-Most of such observations are supported by FLA metrics reported in Table 3.
-However, we note that all four landscapes seem to reveal multimodal structures, as there are dozens to hundreds of local optima with relative large basins (up to a mean size of $2,372$).
-Similar observations have been made by Pushak & Hoos, who speculated that the reasons could be FCNet scenarios fall into the over-parameterized regime, which is different from the other scenarios.
-While we also agree with this reasoning, we seek to conduct further analysis on the local optima using the local optima network (LON) (Appendix B.3).
-Figure 13 (left) shows the LON of FCNet on the parkinsons telemonitoring dataset.
-It can be seen that in general, local optima with better performance tend to feature a large basin of attraction and high in-coming strenghts (i.e., in LON context, the total probability for other local optima to reach it).
-This implies that, despite the pressence of many other local optima, the global optima still plays a pivotal role in the connectivity of the network (see Figure 13, right), which is not the worst case that one may encounter (e.g., a global optimum with tiny basin locates at a secluded region that the optimizer can hardly find.)
-Appendix E Landscape Similarity & Datasets Characteristics
+Appendix F Landscape Similarity & Datasets Characteristics
 Figure 15: Spearman correlation between landscape similarity and dataset characteristics.
@@
 How characteristics of datasets could affect the landscape is an interesting problem to explore in more detail.
-In addition, we also hypothesis that the generalization gap of a model can also depend on the dataset, in addition to its HP setting.
+In addition, we also hypothesize that the generalization gap of a model can also depend on the dataset, in addition to its HP setting.
 To investigate these, we conduct additional experiments on the $57$ tabular datasets to analyze the relationship between landscape similarity and dataset characteristics.
-We first find in the left plot of Figure 15 that the similarity between test and training loss landscapes is positively correlated across models.
+We first find in the left plot of Fig.
+15 that the similarity between test and training loss landscapes is positively correlated across models.
 It implies that on certain datasets, all the $4$ models are more prone to overfit, whereas the opposite could be true for other datasets.
@@
 To further explore which properties of the dataset could potentially contribute to this, we investigated the correlation between train-test landscape similarity and the number of instances & features of each dataset, and their product.
-The results (also in Figure 15 (left)) indicate that the degree of overfitting is correlated with all $3$ dataset size measures.
+The results (also in Fig.
+15 (left)) indicate that the degree of overfitting is correlated with all $3$ dataset size measures.
 In particular, for most models, it would be more likely to encounter overfitting on larger datasets.
 We then proceed to investigate whether correlations between HP loss landscapes induced on different datasets are related to the relative size of the datasets.
-From Figure 15 (right), it is clear to see that pairwise landscape similarities are obviously correlated across models, implying that all the models are likely to induce very different (or the opposite) landscapes on certain pairs of datasets.
+From Fig.
+15 (right), it is clear to see that pairwise landscape similarities are obviously correlated across models, implying that all the models are likely to induce very different (or the opposite) landscapes on certain pairs of datasets.
 We could also see that these pairwise similarities are again correlated with the differences in dataset sizes.
@@
 First, the choice of datasets is only a partial factor contributing to landscape similarity, many other factors like model HPs, training settings can also important roles here.
-Second, the dataset meta-features we used here are rather naive, and there could be more comprehensive features for characterizing dataset properties, e.g., Feurer et al.
-(2015b) had used a collection of $46$ features to assess dataset similarity.
+Second, the dataset meta-features we used here are rather naive, and there could be more comprehensive features for characterizing dataset properties, e.g., [53] had used a collection of $46$ features to assess dataset similarity.
 However, this is beyond the scope of this work and we leave it to future works.
-Appendix F Details of the Experimental Setup
-F.1 Search Spaces
+Appendix G Details of the Experimental Setup
+G.1 Search Spaces
 In this subsection, we elaborate the principles that we follow in designing the search spaces for each model.
-We also provide the detailed hyperparameter grid space for them in Table 4 (XGBoost), Table 5 (DT), Table 6 (RF), Table 7 (RF), Table 8 (CNN) and Table 9 (FCNet).
+We also provide the detailed hyperparameter grid space for them in Table 2 (XGBoost), Table 3 (DT), Table 4 (RF), Table 5 (RF), Table 6 (SVM), and Table 7 (CNN).
 The top-level principle we follow is to include commonly used HPs and exclude unimportant ones, while keeping a good balance between search space coverage and computational cost.
 We first determine the list of HPs to be considered, and to this end:
-  • We surveyed commonly used HPs in the practice (e.g., van Rijn & Hutter identified important HPs of several models using large-scale meta-data from OpenML) and HPO literature (e.g., the search spaces used in HPOBench (Eggensperger et al., 2021)).
-For CNN, we also refer to the design of NAS search spaces (Chitty-Venkata et al., 2023).
+  • We surveyed commonly used HPs in the practice (e.g., [87] identified important HPs of several models using large-scale meta-data from OpenML) and HPO literature (e.g., the search spaces used in HPOBench [88]).
+For CNN, we also refer to the design of NAS search spaces [89].
   • We also run preliminary experiments by fixing all except but one HP to their default values, and vary a single HP using a wide range of values.
@@
 num_leaves $[10,20,30,40,50,100]$ $6$
+Table 6: SVM HP Grid Space ($2340$ Configurations)
+Hyperparameter Grid Values Count
+gamma $[1\mathrm{e}^{-4},10]$ $45$
+c $[1\mathrm{e}^{-3},\dots,1\mathrm{e}^{3}]$ $54$
 Table 7: CNN HP Grid Space ($6,480$ Configurations)
@@
 learning_rate [$1\mathrm{e}^{-4},5\mathrm{e}^{-4},1\mathrm{e}^{-3},3\mathrm{e}^{-3},5\mathrm{e}^{-3}$] $5$
-batch_size $[128,256]$ $10$
+batch_size [$128,256$] $10$
 drop_out [$0.0,0.1,0.25,0.5$] $4$
-width_linear $[256,512,1024]$ $3$
-width_conv $[32,64,128]$ $3$
-#conv_block $[2,4,6]$ $3$
-Table 9: FCNet HP Grid Space ($62,208$ Configurations)
- Hyperparameter Grid Values Count
-initial_LR $\{5\mathrm{e}^{-4},1\mathrm{e}^{-3},5\mathrm{e}^{-3},1\mathrm{e}^{-2},1\mathrm{e}^{-2},.1\mathrm{e}^{-1}\}$ $6$
-batch_size $\{8,16,32,64\}$ $4$
-LR_schedule {cosine, fix} $2$
-activation_layer_1 {relu, tanh} $2$
-activation_layer_2 {relu, tanh} $2$
-layer_1_Size $\{16,32,64,128,256,512\}$ $6$
-layer_2_Size $\{16,32,64,128,256,512\}$ $6$
-dropout_Layer_1 $\{0.0,0.3,0.6\}$ $3$
-dropout_Layer_2 $\{0.0,0.3,0.6\}$ $3$
-F.2 Datasets
+width_linear [$256,512,1024$] $3$
+width_conv [$32,64,128$] $3$
+#conv_block [$2,4,6$] $3$
+G.2 Datasets
 Here we provide basic information regarding the $5$ groups of datasets used for this study, including: $i)$: numerical regression (Table 8), $ii)$: numerical classification (Table 9), $iii)$: categorical regression (Table 10), $iv)$: categorical classification (Table 9), and $v)$: image classifcation (Table 12).